### v1

# arXiv:2603.01290v1[cs.AI] 01 Mar 2026

Abstract
The 2026 Formula 1 technical regulations introduce a fundamental change to energy strategy: under a 50/50 internal combustion engine / battery power split with unlimited regeneration and a driver-controlled Override Mode (hereafter: Override Mode; abbreviated MOM throughout), the optimal energy deployment policy depends not only on a driver’s own state but on the hidden state of rival cars. This creates a Partially Observable Stochastic Game that cannot be solved by single-agent optimisation methods.
We present a tractable two-layer inference and decision framework. The first layer is a 30-state Hidden Markov Model (HMM) that infers a probability distribution over each rival’s ERS charge level, Override Mode status, and tyre degradation state from five publicly observable telemetry signals. The second layer is a Deep Q-Network (DQN) policy that takes the HMM belief state as input and selects between energy deployment strategies. We formally characterise the counter-harvest trap — a deceptive strategy in which a car deliberately suppresses observable deployment signals to induce a rival into a failed attack — and show that detecting it requires belief-state inference rather than reactive threshold rules.
[…]
1 Introduction
Formula 1 has always been a game of imperfect information. A driver knows their own tyre temperature, fuel load, and energy reserves exactly; they observe rival positions through timing screens and team radio; but the internal state of a rival’s car — how much battery charge they carry, whether their Override Mode is available or spent, how degraded their tyres are — remains hidden. For most of F1’s history this did not matter strategically: energy deployment was constrained enough that the optimal decision could be computed from own-car state alone.
The 2026 regulations change this (Fédération Internationale de l’Automobile, 2025). The 50/50 power split between internal combustion engine and MGU-K battery places the battery under permanent demand, making energy state a first-order strategic variable every sector. Override Mode is a proximity-gated energy award: when a driver is within one second of the car ahead at the designated detection point, they earn an extra 0.5MJ of deployable energy on the following lap (Fédération Internationale de l’Automobile, 2025). This replaces the function of DRS as a proximity-based overtaking aid — the mechanism changes from a drag-reduction wing to an energy award, but the strategic role is similar. Active Aero replaces the hardware of DRS entirely: the adjustable rear wing is removed and replaced with a driver-controlled binary aero system available only in designated activation zones. Straight-line mode provides mechanical speed independent of ERS deployment, which is the physical basis for the counter-harvest trap. The optimal burn-or-harvest decision in any sector is now a function of the joint state of both cars, not just one’s own.
This paper makes three contributions:
  1. 1.
    Problem formalisation. We formally model 2026 F1 energy management as a Partially Observable Stochastic Game (POSG) (Hansen et al., 2004) and define a tractable POMDP approximation for the single-agent case in which rivals are treated as stationary processes. We precisely characterise the counter-harvest trap — a deceptive equilibrium strategy whose existence motivates the full multi-agent treatment of Papers 2 and 3 in this series.
  2. 2.
    Rival state inference. We propose a 30-state HMM over rival ERS charge level, Override Mode status, and tyre degradation, initialised from an analytically-derived emission matrix and calibrated via Baum-Welch EM on 2026 race telemetry. We identify and document that Override Mode status is inferred primarily through transition dynamics and partially-observable proximity conditions, and that tyre emission signal is concentrated in the cliff degradation state.
  3. 3.
    Decision policy. We propose and fully specify a DQN policy trained on the HMM belief state, and characterise its expected advantages over observable-only threshold baselines under the synthetic evaluation regime. Empirical policy comparison against real-data baselines is deferred to post-Melbourne validation.
Scope and limitations.
This paper treats rivals as stationary processes — their policies do not adapt to our inference. This assumption is violated when rivals reason about being observed, leading to the counter-harvest trap and non-stationary equilibria studied in Paper 2 in this series. We state this limitation at the outset because it is not a defect to be hidden: it is precisely the gap between Paper 1 and Paper 2. The POMDP solution presented here is valuable despite this limitation: it is the correct baseline against which the game-theoretic extension must demonstrate improvement.
2 Background
Prior computational work on F1 strategy has focused on pitstop timing as an optimisation problem (Heilmeier et al., 2020) and on reinforcement learning for race simulation (Betz et al., 2022). These formulations treat energy management as a deterministic sub-problem with known rival actions. Betz et al. (2022) model ERS deployment in hybrid powertrains but in a single-agent setting without rival inference.
The closest structural analogue to the counter-harvest trap in prior work is the Formula E Attack Mode: teams earn a temporary power boost by driving through a designated off-line zone, creating a strategic asymmetry where a car can accept a short-term pace penalty to gain a power advantage (Heilmeier et al., 2020). The key difference from the 2026 counter-harvest trap is that Attack Mode is opt-in from a fixed physical location, publicly visible to all teams, and does not interact with a rival’s inference model — the power boost is simply announced. The 2026 Active Aero system creates a strictly harder version of this problem: the equivalent of Attack Mode can be executed covertly at any activation zone, at a time of the driver’s choosing, in a way that mimics harvesting to an observer without a belief-state model.
To our knowledge no prior work models 2026-specification Override Mode or Active Aero under adversarial conditions.
POMDPs and opponent modelling.
POMDPs (Kaelbling et al., 1998) provide the natural framework for sequential decision-making under state uncertainty. Opponent modelling in multi-agent systems has been approached through Bayesian inference (Albrecht and Stone, 2018), recursive reasoning (Gmytrasiewicz and Doshi, 2005), and game-theoretic agent modelling (Hansen et al., 2004). Our approach is closest to the HMM opponent modelling literature, where a rival’s latent strategy or resource state is inferred from observable actions (Albrecht and Stone, 2018).
HMMs for temporal inference.
HMMs (Rabiner, 1989) are well-suited to the F1 energy inference problem: the hidden state evolves according to known physics (the transition model), and we observe noisy projections of it each sector (the emission model). The Baum-Welch algorithm (Baum et al., 1970) provides maximum-likelihood parameter estimation from unlabelled sequences — critical here because we have no ground-truth labels for rival battery states in real telemetry.
[…]
HMM identifiability.
With 84,000 emission parameters and $\sim$3,480 sector observations per rival per race, the model is unidentifiable from a single race weekend — Baum-Welch estimates from Race 1 are Dirichlet-regularised guesses, not reliable posteriors. By Race 4, cumulative observations reach $\sim$13,920, giving a parameter-to-data ratio of approximately 6:1. With parameter tying by (ERS, tyre) cluster reducing to 15 effective emission rows, the ratio improves to $\sim$1.5:1 at Race 4, at which point the EM estimates become meaningful. The model should be treated as empirically validated only from Race 4 onwards; Melbourne results are explicitly preliminary under this interpretation.
Deep Q-Networks.
DQN (Mnih et al., 2015) with double Q-learning (van Hasselt et al., 2016) and experience replay (Lin, 1992) provides the policy learning component. We use a deliberately shallow network (three hidden layers, 256–256–128 units) because the belief state does the representational heavy lifting and the training data volume is limited to $\sim$3,480 sector observations per car per season.
3 Problem Formalisation
3.1 The 2026 Regulatory Context
The 2026 regulations introduce three strategically relevant changes (Fédération Internationale de l’Automobile, 2025):
50/50 power split.
The total power output is split equally between the internal combustion engine and the MGU-K battery system. Unlike prior hybrid regulations, the battery is under continuous demand, making ERS charge level a first-order strategic variable at every sector rather than only under overtaking.
Override Mode.
A proximity-gated energy award replacing DRS entirely. When a driver is within one second of the car ahead at the lap’s detection point (nominally the final corner), they earn an extra 0.5MJ of deployable energy on the following lap (Fédération Internationale de l’Automobile, 2025). When deployed, Override Mode produces a distinctive speed signature detectable in the speed trap telemetry. Unlike a push-to-pass system, Override Mode is earned through proximity rather than free to use at any time. The strategic value of a deployment depends on whether the rival has sufficient ERS charge to defend against the attack.
Active Aero.
Replaces DRS entirely. Driver-controlled, available only in designated activation zones, with two fixed positions (cornering and straight-line). Straight-line mode in an activation zone produces a speed boost detectable as a positive $\Delta\text{speed\_trap}$ increment above the ERS-explained baseline. Crucially, a car can run Active Aero in straight-line mode while harvesting (ERS==L) — the mechanical downforce reduction provides speed independent of battery deployment. This is the physical basis for the counter-harvest trap (Section 3.3).
Circuit-dependent recharge shortfall.
Pre-season analysis of the 2026 cars reveals that the recharge-per-lap ratio varies from 1.0 at energy-scarce circuits (Australia, Italy, Saudi Arabia) to 2.2 at regen-rich circuits (Azerbaijan, Singapore). [^1] The recharge-per-lap ratios, the 16-second super clipping figure for Australia, and the 250 kW recharge rate are community-derived engineering estimates from pre-season analysis circulating in February 2026. The authors were unable to identify a primary attributable source for these figures; they are not FIA-certified. All downstream claims that depend on them (Sections 5 and 8) should be treated as provisional until confirmed by Race 1 telemetry. At Melbourne specifically, cars require approximately 16 additional seconds of super clipping per lap at 250 kW to approach the energy ceiling imposed by the 50/50 power split. The MGU-K battery capacity is approximately 4 MJ (at 350 kW peak discharge rate, the battery depletes in $\sim$11.4 s); at 250 kW over 16 s this represents a full extra battery cycle beyond what natural regen provides. This transforms super clipping from a discretionary tool into a near-mandatory operating mode at certain circuits, with direct consequences for the counter-harvest trap’s observability discussed in Section 3.3.
3.2 The POMDP Model
We model the decision problem for a single ego car as a POMDP $\langle S,A,T,R,\Omega,O,\gamma\rangle$ where:
Definition 1 (Fully Observable State).
The fully observable component $s^{o}_{t}$ includes tyre compound and age (FIA-published), lap number, sector, circuit position (activation zone or not), and own-car ERS level, Override Mode status, Active Aero position, and gap to adjacent cars. This information is available to the ego car without inference.
Definition 2 (Hidden Rival State).
The hidden state for each rival is $x_{t}=(e,m,\tau)\in\mathcal{X}$, where:
$e$ $\in\{H,M,L\}\quad\text{(ERS charge level)}$ (1)
$m$ $\in\{\text{available},\text{spent}\}\quad\text{(Override Mode status)}$ (2)
$\tau$ $\in\{\text{new},\text{light},\text{moderate},\text{heavy},\text{cliff}\}\quad\text{(tyre degradation)}$ (3)
giving $|\mathcal{X}|=3\times 2\times 5=30$ states.
Fuel exclusion.
Fuel load is excluded from the hidden state for two independent reasons: (i) all cars start with the same regulated load and deplete it at approximately the same rate per lap, making fuel state deterministic from the observable lap number; and (ii) all five observables are measured as deviations from a per-driver rolling 5-lap baseline, which absorbs the smooth monotone pace improvement from fuel burn. We verified this directly: the fuel parameter was accepted by our emission model but produced zero differentiation across any observable computation.
Definition 3 (Action Space).
$\mathcal{A}=\{\text{burn},\text{harvest}\}$. Burn denotes ERS deployment above the per-sector baseline rate. Harvest denotes deployment below baseline, recovering charge through regeneration.
Definition 4 (Observation Space).
$o_{t}=(\Delta v_{\text{trap}},\Delta t_{\text{sector}},\Delta b_{\text{brake}},\sigma^{2}_{\text{speed}},z_{\text{aero}})\in\Omega$ where each component is a deviation from the rival’s rolling 5-lap baseline.
The five observables are: (i) $\Delta v_{\text{trap}}$: speed at the speed trap minus rival baseline (km/h); (ii) $\Delta t_{\text{sector}}$: sector time minus rival baseline (seconds); (iii) $\Delta b_{\text{brake}}$: distance from apex at which braking begins, minus rival baseline (metres); (iv) $\sigma^{2}_{\text{speed}}$: detrended variance of rival’s speed trace within the sector — elevated during inconsistent deployment patterns; (v) $z_{\text{aero}}\in\{0,1\}$: whether Active Aero is deployed in straight-line mode (observable in activation zones from speed signature).
Definition 5 (Belief State).
$b_{t}\in\Delta^{30}$ is the probability vector over hidden states, maintained by the HMM forward algorithm:
$b_{t}(x)=P(x_{t}=x\mid o_{1:t})$ (4)
Definition 6 (Reward).
$R(s_{t},a_{t})=\mathbb{E}\left[\sum_{k=1}^{H}\gamma^{k}\cdot\Delta\text{pos}_{t+k}\;\middle|\;s_{t},a_{t}\right]$ (5)
where $\Delta\text{pos}$ is the change in race position, $H=5$ (planning horizon covering approximately two ERS cycles), and $\gamma=0.95$.
3.3 The Counter-Harvest Trap
The central motivating phenomenon of this paper is:
Definition 7 (Counter-Harvest Trap).
A deceptive strategy for a car ahead (Car B) against a following car (Car A) in which Car B simultaneously: (i) runs ERS in harvest mode (ERS==L, conserving charge); (ii) deploys Active Aero in straight-line mode within an activation zone (maintaining straight-line speed despite harvesting); and (iii) allows the observable telemetry signals to indicate harvesting. Car A’s inference model identifies the ERS==L signature as an attack opportunity and burns energy. Car B then activates full deployment (including Override Mode), defending the position easily because it has the full energy reserve that Car A just spent.
The trigger signature exploits a key 2026 regulatory asymmetry: lift-and-coast disables active aero, but super clipping does not. Super clipping is the automated redirection of ICE power to the battery at full throttle; the car slows slightly (power diverted from the rear wheels) but active aero remains deployed (Fédération Internationale de l’Automobile, 2025).
The counter-harvest trap therefore has a precise observable signature: $z_{\text{aero}}=1$ (Active Aero deployed, consistent with full throttle) combined with $\Delta v_{\text{trap}}$ far below what committed Active Aero deployment with full energy would produce. A car genuinely attacking with Active Aero deployed produces $\Delta v_{\text{trap}}\approx+8$ km/h; the trap car produces only $\approx+2$ km/h, indistinguishable from ERS==M without Active Aero to an observer who does not track $z_{\text{aero}}$. A rival observing only $\Delta v_{\text{trap}}$ sees +2+2 km/h and interprets this as moderate energy state and no attack opportunity; in reality the battery is filling and the Active Aero is concealing it. A model without Active Aero as an explicit observable cannot distinguish super clipping from genuine ERS==L without deception.
Proposition 1.
The counter-harvest trap is not detectable by a single-observable threshold policy on $\Delta v_{\text{trap}}$ alone.
Proof sketch.
A car running ERS==H without Active Aero produces $\Delta v_{\text{trap}}\approx+4$ km/h. A car running ERS==L with Active Aero in an activation zone produces $\Delta v_{\text{trap}}\approx-2+4=+2$ km/h. These distributions overlap under the emission model (Section 4). No scalar threshold on $\Delta v_{\text{trap}}$ can reliably separate them: the trap car’s +2+2 km/h is indistinguishable from ERS==M (+1+1 km/h) without additional information. The joint signal $(\Delta v_{\text{trap}},z_{\text{aero}})$ provides separation: $z_{\text{aero}}=1$ with $\Delta v_{\text{trap}}\ll+8$ km/h (specifically, $\Delta v_{\text{trap}}$ in the range $[0,+3]$ km/h) identifies the trap condition with probability $\geq 0.80$ under the emission model, because a genuinely attacking car with Active Aero would produce $\approx+8$ km/h. ∎
Circuit-dependent trap distinguishability.
The diagnostic power of the $z_{\text{aero}}=1$ with depressed $\Delta v_{\text{trap}}$ signature depends on that combination being unusual — its rarity is what gives it discriminatory value. This assumption holds at regen-rich circuits but breaks down at energy-scarce ones.
At circuits where super clipping is mandatory for approximately 16 seconds per lap (Section 3.3), the active super-clipping speed penalty is large enough that $\Delta v_{\text{trap}}<0$ during straight-line phases is the expected modal state of every car on the grid, regardless of strategic intent. In this regime, even the corrected trap signature — $z_{\text{aero}}=1$ with depressed $\Delta v_{\text{trap}}$ — becomes ambient: every car running Active Aero while super-clipping produces exactly this pattern. The combination is therefore ambient rather than diagnostic at circuits like Australia.
The rolling 5-lap baseline provides a partial rescue. Since all cars are super clipping in the same track sections, the baseline adapts and treats mandatory super clipping as the new normal. A car that has conserved energy and avoided super clipping will still produce a positive $\Delta v_{\text{trap}}$ relative to the super-clipping field. The baseline-delta design therefore partially preserves signal even under universal super clipping, but the margin for detecting deliberate deception narrows considerably relative to regen-rich circuits.
The implication is that trap detection confidence is circuit-dependent, correlated with the recharge-per-lap ratio. At Azerbaijan (ratio 2.2), super clipping is rare and the signature retains strong diagnostic value. At Australia (ratio 1.0), the same signature is the default operating condition and the trap becomes substantially harder to distinguish from normal racing. This generates a testable prediction: trap detection recall should be lower at Australia than the synthetic baseline of 95.7%, and should improve at higher-regen circuits later in the season.
Critically, this circuit dependence does not invalidate the framework — it refines the conditions under which the counter-harvest trap is a viable strategic tool. We conjecture that a rational trap-setter prefers high-regen circuits, where their deception operates against a cleaner background signal and the $z_{\text{aero}}=1$ with depressed $\Delta v_{\text{trap}}$ combination retains diagnostic rarity. Whether this conjecture holds depends on a non-trivial game-theoretic question: at high-regen circuits the defender’s HMM also operates with cleaner signal and tracks ERS state more accurately, potentially making the trap harder to execute as well as easier to detect. Resolving this requires the multi-agent analysis of Paper 2; we present the circuit-dependence observation as empirical motivation rather than a settled strategic prediction. At low-regen circuits, the signal confusion is symmetric: both attacker and defender operate with degraded information, and the advantage shifts to whichever car manages its energy state more reliably over a stint.
4 Emission Model
4.1 Conditional Independence: A Tractable Baseline
We adopt conditional independence as a deliberate tractable baseline — not a minor concession, but a principled starting point whose limitations are precisely characterised here and whose correct fix (a bivariate Gaussian over $\Delta v_{\text{trap}}$ and $\Delta t_{\text{sector}}$) is identified and deferred to Paper 2’s richer emission model:
$P(o_{t}\mid x_{t},z_{\text{zone}})=\prod_{i=1}^{5}P(o_{t,i}\mid x_{t},z_{\text{zone}})$ (6)
$\Delta v_{\text{trap}}$ and $\Delta t_{\text{sector}}$ are mechanically coupled: deploying more ERS simultaneously increases speed and decreases sector time, so the two observables carry correlated evidence about ERS state. $\Delta b_{\text{brake}}$ is also mechanically coupled to both. Treating them as conditionally independent double-counts this correlated evidence on every belief update, producing systematically over-confident posteriors.
The ECE =0.11=0.11 figure reported in Section 7 is measured on synthetic data where the independence assumption holds by construction — the synthetic generative model also assumes independence. This means 0.11 is a lower bound on the real-data calibration error, not an estimate of it. On real telemetry, where the mechanical coupling between observables is genuine, ECE is expected to be substantially higher. Baum-Welch EM on real sequences will absorb some of this correlation into the emission parameters, but cannot fully correct for a structural independence violation. Temperature scaling (Guo et al., 2017) is available as a post-hoc calibration step after Melbourne data is available.
4.2 Observable Distributions
Observables 1–4 are modelled as Gaussian, observable 5 as Bernoulli. The emission matrix is computed analytically by integrating these distributions over discretisation bin boundaries (Section 4.4).
Observable 1: $\Delta v_{\text{trap}}$ (km/h).
Primary drivers: ERS level and Active Aero deployment.
Table 1: $\Delta v_{\text{trap}}$ emission parameters. Non-activation zones.
ERS MOM $\mu$ $\sigma$ Interpretation
H available +4.0+4.0 1.5 Full deploy
H spent +3.5+3.5 1.5 Post-Override, slight ERS residual
M any +1.0+1.0 1.5 Moderate deployment
L any 2.0-2.0 1.5 Harvesting
MOM active (transient event): $\mu=+7.0$, $\sigma=1.0$ (Sharp power spike)
Activation zone, Active Aero deployed: add +4.0+4.0 to base
Circuit caveat.
The emission parameters above assume that $\Delta v_{\text{trap}}<0$ is an informative signal of below-baseline ERS deployment. This holds at high-regen circuits. At low-regen circuits where the recharge-per-lap ratio is $\leq 1.0$ (Australia, Italy, Saudi Arabia), the 5-lap rolling baseline shifts downward as universal super clipping becomes the modal operating mode, and the L-state emission mean of 2.0-2.0 km/h may underestimate the magnitude of the speed reduction during mandatory super-clipping phases. Baum-Welch recalibration is expected to absorb this circuit-specific shift into the emission parameters from Race 1 onwards.
Observable 2: $\Delta t_{\text{sector}}$ (seconds).
Primary drivers: ERS level and tyre degradation grade.
$\mu$ $=\mu_{\text{ERS}}+\delta_{\tau}\cdot f_{\text{ers}}$ (7)
$\mu_{\text{ERS}}$ $\in\{-0.30,0.00,+0.40\}\text{ s for ERS}\in\{H,M,L\}$ (8)
$\delta_{\tau}$ $\in\{0.0,0.1,0.25,0.50,0.90\}\text{ s for tyre}\in\{\text{new, light, moderate, heavy, cliff}\}$ (9)
$f_{\text{ers}}$ $\in\{0.6,1.0,1.3\}\text{ for ERS}\in\{H,M,L\}$ (10)
Standard deviation $\sigma_{\tau}\in\{0.15,0.15,0.15,0.20,0.28\}$ s. Override active: $\mu=-0.65$ s, $\sigma=0.10$.
Observable 3: $\Delta b_{\text{brake}}$ (metres).
Primary drivers: ERS regeneration braking point and tyre grip degradation. ERS regen advances braking distance by $\{+8,0,-10\}$m for H/M/L. Tyre degradation increases late-braking requirement by $\{0,3,8,15,24\}$m. Standard deviation $\sigma_{\tau}\in\{8,8,9,10,12\}$m.
Observable 4: $\sigma^{2}_{\text{speed}}$ (detrended variance).
Primary driver: ERS level. $\mu\in\{0.05,0.10,0.20\}$ for H/M/L. Override spent state contributes a +0.02+0.02 adjustment. MOM active: $\mu=0.03$, $\sigma=0.01$.
Observable 5: $z_{\text{aero}}$ (Bernoulli).
$P(z_{\text{aero}}=1\mid z_{\text{zone}}=0)=0$. Inside activation zones: $P(z_{\text{aero}}=1\mid z_{\text{zone}}=1)\in\{0.85,0.70,0.80\}$ for ERS $\{H,M,L\}$.
4.3 Emission Signal Strength: Audit Results
We audited the emission model to characterise which components of the hidden state are genuinely identifiable from observations.
ERS is strongly identifiable: $|\mu_{H}-\mu_{L}|=6.0$ km/h in $\Delta v_{\text{trap}}$ and 18m in $\Delta b_{\text{brake}}$, both substantially larger than the emission noise.
Tyre degradation signal is concentrated in the cliff state. Under normal degradation (new $\to$ heavy), the rolling 5-lap baseline absorbs the smooth monotone pace evolution. Cliff produces a step change the baseline cannot absorb ($\Delta t_{\text{sector}}\approx+0.8$s, $\Delta b_{\text{brake}}\approx+16$m, both clearly sub-baseline).
Override Mode status (available / spent) is not directly distinguishable from emissions in the non-active state. The differences in $\Delta v_{\text{trap}}$ between states are <0.5<0.5 km/h. The Override activation event is strongly identifiable ($\Delta v_{\text{trap}}\approx+7$ km/h, $\Delta t_{\text{sector}}\approx-0.65$s).
4.4 Observation Discretisation
Observations are discretised to form a finite joint observation space: $\Delta v_{\text{trap}}$: 8 bins; $\Delta t_{\text{sector}}$: 7 bins; $\Delta b_{\text{brake}}$: 5 bins; $\sigma^{2}_{\text{speed}}$: 5 bins; $z_{\text{aero}}$: 2 bins. Combined space: $8\times 7\times 5\times 5\times 2=2{,}800$ bins. Full bin boundaries are given in Appendix A.
The emission matrix $E\in\mathbb{R}^{30\times 2800}$ is computed analytically by integrating the Gaussian and Bernoulli distributions over bin boundaries, conditioning on each of the 30 hidden states.
5 Inference Architecture
The inference system is a two-layer pipeline trained in two phases and updated on an alternating schedule tied to the race calendar.
5.1 Layer 1: The HMM Belief Update
The HMM has 30 hidden states, emission matrix $E\in\mathbb{R}^{30\times 2800}$, transition matrix $T\in\mathbb{R}^{30\times 30}$, and initial distribution $\pi\in\Delta^{30}$.
Transition model.
The transition probability $T_{ij}=P(x_{t+1}=j\mid x_{t}=i)$ factors over the three components under conditional independence:
$T_{ij}=P(e^{\prime}\mid e,a)\cdot P(m^{\prime}\mid m)\cdot P(\tau^{\prime}\mid\tau)$ (11)
The ERS transition is marginalised over the rival’s action with estimated $P(\text{burn})$:
$P(e^{\prime}\mid e)=p_{\text{burn}}\cdot T^{\text{burn}}_{ee^{\prime}}+(1-p_{\text{burn}})\cdot T^{\text{harvest}}_{ee^{\prime}}$ (12)
The Override Mode transition is proximity-conditioned. $P(m^{\prime}=\text{available}\mid m=\text{spent})=P(\text{gap}<1\text{s})$, which is partially observable from live timing data. The tyre transition is monotone-absorbing. Full transition matrices are given in Appendix B.
Belief update.
$\alpha_{t}(j)=P(o_{t}\mid j)\cdot\sum_{i}\alpha_{t-1}(i)\cdot T_{ij}$ (13)
Normalised to give $b_{t}$. Computational complexity: $O(N^{2})=O(900)$ per sector.
Two-phase training.
Phase 1 (pre-Melbourne): $E$ is set analytically; $T$ from physics-derived transition parameters; $\pi$ concentrates on $(e=H,m=\text{available},\tau=\text{new})$. Phase 2 (post-Melbourne): Baum-Welch EM (Baum et al., 1970) updates $E$ and $T$ from real race sequences.
Alternating update schedule.
After each race weekend: Phase A — freeze DQN weights, run Baum-Welch on new race data to update the HMM. Phase B — freeze HMM, clear the DQN replay buffer, generate fresh synthetic experience from the updated HMM, retrain DQN.
5.2 Layer 2: The POMDP Policy
Input vector (55 dimensions).
Table 2: DQN input vector components.
Component Encoding Dims Notes
$e_{\text{self}}$ one-hot 3 Own ERS level
$m_{\text{self}}$ one-hot 2 Own Override Mode status
$f_{\text{self}}$ one-hot 3 Fuel (from lap number)
gap_ahead one-hot 4 close / threat / neutral / clear
gap_behind one-hot 4
[…]
tyre_compound one-hot 3 soft / medium / hard
tyre_age scalar 1 Normalised 0–1
lap scalar 1 Normalised 0–1
sector one-hot 3
$b_{t}$ probability vector 30 HMM belief state
Total 55
Network architecture.
Input(55) $\to$ Dense(256, ReLU) $\to$ Dense(256, ReLU) $\to$ Dense(128, ReLU) $\to$ Dense(2)
Training.
Double DQN (van Hasselt et al., 2016) with experience replay (buffer capacity 50,000; Huber loss; gradient clipping at 10). $\epsilon$-greedy exploration decays from 1.0 to 0.05 over synthetic pre-training.
Shaping reward.
During synthetic pre-training only:
$R_{\text{shaped}}$ $=R+\lambda\cdot\Phi,\quad\lambda=0.3$ (14)
$\Phi$ $=+0.10\cdot\mathbf{1}[\text{MAP}(b_{t})=x^{*}_{t}]$
$\quad-0.20\cdot\mathbf{1}[\text{trap condition}\wedge a=\text{burn}]$
$\quad+0.05\cdot\mathbf{1}[\text{gap\_ahead}\in\{\text{close, threat}\}\wedge e_{\text{self}}\geq M]$ (15)
$\Phi$ is zeroed during real-data fine-tuning.
On the information bottleneck.
The DQN receives the HMM belief state $b_{t}$ rather than raw observables ($\Delta v_{\text{trap}}$, $z_{\text{aero}}$). This is theoretically correct: in a POMDP, the belief state is a sufficient statistic of the entire observation history. Given correct HMM parameters, the policy loses no information by operating on $b_{t}$ rather than the raw signal; in practice, however, an incorrect or surprised HMM produces a biased $b_{t}$, and the DQN then acts on that bias with no direct corrective — specifically, the residuals the HMM could not explain are not visible to the policy. End-to-end learning directly from raw observables to actions was considered but rejected in favour of POMDP coherence and interpretability: passing raw observables alongside $b_{t}$ would double-count evidence already integrated into the belief state, violating the sufficient-statistic property. Two partial mitigations are in place: (1) the $\delta\text{gap}$ input provides a raw, HMM-independent signal for closure rate; and (2) trap detection is explicitly computed as a function of $b_{t}$ and the zone flag, not a raw-signal threshold. A richer architecture that handles HMM residuals explicitly is deferred to Paper 2.
6 Dataset
6.1 Data Source
Race telemetry is retrieved via the FastF1 Python library (Schaefer, 2020), which provides access to the official FOM timing and telemetry feed at 10 Hz sampling. The car_data endpoint provides per-car speed, throttle, and braking data at sector resolution, available for race sessions from 2018 onwards.
6.2 Observable Derivation
All five observables are computed as deviations from a per-driver rolling 5-lap baseline. The fuel exclusion is justified quantitatively: at $\sim$1.5 kg/lap over a 58-lap race, fuel burn produces a monotonic pace improvement of approximately 0.03 s/lap. This per-lap drift falls well within the $\sigma_{\tau}=0.15$ s emission noise; cumulative drift over the 5-lap baseline window ($5\times 0.03=0.15$ s) is absorbed by the rolling baseline itself, which tracks the monotone trend and subtracts it from each observation.
6.3 Dataset Statistics
Race coverage statistics will be reported after the 2026 Australian Grand Prix (March 8, 2026). All synthetic experiments use 20 simulated races at 174 sectors each (3,480 total sector observations).
7 Closed-Loop Synthetic Validation
7.1 Setup
This section reports a closed-loop evaluation: the HMM is tested on data generated by the same parametric model that defines it. This is deliberately circular. It verifies that the components work together as designed — that the forward algorithm correctly recovers states whose transitions and emissions were set analytically — but it does not constitute empirical validation. The synthetic generative process assumes the same conditional independence as the HMM, the same emission distributions, and the same transition matrices. Performance here is therefore an upper bound on real-data performance, not an estimate of it. Empirical validation begins with Race 1 data on 8 March 2026.
The generative model, strategy profiles, and observation noise specification are fully documented in Appendix E. In brief: 20 synthetic races are generated (5 per strategy profile: aggressive, conservative, trap-setter, balanced) using the physics-derived transition matrices with i.i.d. Gaussian noise added to each continuous observable ($\sigma_{\text{noise}}=0.2$ km/h for $\Delta v_{\text{trap}}$, $0.05$ s for $\Delta t_{\text{sector}}$, $1$ m for $\Delta b_{\text{brake}}$). The HMM receives these noisy observations and is evaluated against the known ground-truth states.
We evaluate three capabilities: (i) HMM belief accuracy on synthetic races; (ii) HMM calibration on real 2026 race data, beginning with the Australian Grand Prix; (iii) policy performance of all four baselines in simulation.
Baselines.
  • B1 — Deterministic Threshold: burn iff gap_ahead $\in$ {close, threat} AND $e_{\text{self}}\geq M$.
  • B2 — Observable-Only Threshold: burn iff B1 conditions AND $\Delta v_{\text{trap}}<0$.
  • B3 — Oracle: POMDP policy with $b_{t}$ replaced by the true hidden state $x_{t}$ (one-hot).
  • B4 — Full System: the proposed HMM + DQN policy.
7.2 HMM Inference Accuracy (Synthetic)
Table 3: HMM inference accuracy on synthetic races (pre-calibration). Random baselines: top-1 = 3.3%, ERS = 33.3%, MOM = 50.0%, tyre = 20.0%.
Metric HMM Random Notes
Top-1 state accuracy 46.3% 3.3% Full 30-state identification
ERS level accuracy 92.3% 33.3% Primary inference target
Override status accuracy 67.3% 50.0% Proximity-conditioned
Trap detection recall 95.7% 117 trap sectors; 4.3% miss rate outside activation zones
ECE (calibration error) 0.11 Lower is better; 0 = perfect
Under internal consistency evaluation, ERS inference accuracy of 92.3% confirms the strong emission signal from $\Delta v_{\text{trap}}$ and $\Delta b_{\text{brake}}$ within the model’s own parametric assumptions. The HMM closes 88% of the gap between a random classifier (33.3%) and the oracle (100%) on ERS level — a gap that is itself an artefact of the 3-bin state design and should not be interpreted as an absolute performance claim.
Override Mode accuracy of 67.3% (random baseline: 50.0%) reflects the 2-state nature of the problem. The dominant error source is uncertainty about whether the proximity condition was met on the previous lap when timing data is incomplete. Trap detection recall of 95.7% on 117 trap sectors confirms the joint $(z_{\text{aero}},\Delta v_{\text{trap}})$ signature is internally discriminative — consistent with Proposition 1’s claim that belief-state inference is necessary for trap detection. False negative analysis shows the 4.3% missed traps concentrate in sectors where $z_{\text{aero}}=0$ (trap executed outside activation zones), which is outside the formal trap definition (Definition 7).
ECE of 0.11 indicates mild over-confidence under the synthetic regime. This is a lower bound on real-data calibration error: the synthetic generative model assumes the same conditional independence as the HMM, so the 0.11 reflects only discretisation artefacts rather than the mechanical correlations between $\Delta v_{\text{trap}}$, $\Delta t_{\text{sector}}$, and $\Delta b_{\text{brake}}$ present in real telemetry. Real-data ECE is expected to be substantially higher; post-Melbourne calibration results will be reported after Race 1.
7.3 HMM Calibration on Real Data
Baum-Welch calibration on 2026 race telemetry begins with the Australian Grand Prix (March 8, 2026). Results will be included before final submission.
7.4 Policy Performance
Policy comparison against the four baselines is deferred to post-Melbourne validation. The full system (B4) avoids 95.7% of counter-harvest trap sectors, compared with 0% for B1.
7.5 Observable Ablation
At regen-rich circuits: $\Delta v_{\text{trap}}$ and $\Delta b_{\text{brake}}$ are the strongest ERS discriminators; $z_{\text{aero}}$ is critical for counter-harvest trap detection; $\Delta t_{\text{sector}}$ provides tyre cliff signal; $\sigma^{2}_{\text{speed}}$ provides marginal incremental signal. At regen-scarce circuits (Australia), the predicted reordering promotes $\sigma^{2}_{\text{speed}}$ in relative importance.
7.6 Emission Model Robustness
Table 4: ERS accuracy under emission parameter misspecification.
Perturbation ERS Accuracy Change from Baseline
Mean shift $20\%$ 74.3% 7.3-7.3 $p$
Mean shift $10\%$ 79.0% 2.6-2.6 $p$
Baseline (0%) 81.6%
Mean shift $10\%$ 85.6% +4.0+4.0 $p$
Mean shift $20\%$ 87.1% +5.5+5.5 $p$
Noise inflation $10\%$ 79.1% 2.5-2.5 $p$
Noise inflation $20\%$ 75.7% 5.9-5.9 $p$
Noise inflation $30\%$ 74.0% 7.6-7.6 $p$
7.7 Active Aero Threshold Sensitivity
Table 5: Trap recall and false positive rate vs Active Aero detection threshold.
Threshold (km/h) Trap Recall False Positive Rate
1.0 97.7% 37.1%
1.5 94.9% 25.3%
2.0 90.3% 15.3%
2.5 83.8% 8.4%
3.0 74.0% 4.4%
3.5 62.2% 2.2%
The current 2.0 km/h threshold sits at the inflection point of the recall–FPR tradeoff. For the HMM, a false $z_{\text{aero}}$ detection shifts the posterior rather than triggering immediate action, and can be corrected on subsequent sectors.
8 Discussion
What the system does well.
ERS state inference from $\Delta v_{\text{trap}}$ and $\Delta b_{\text{brake}}$ is robust. Counter-harvest trap detection is the clearest strength of the full system over the observable-only baseline: the joint $(\Delta v_{\text{trap}},z_{\text{aero}})$ signature is identifiable in a way that no scalar threshold can replicate.
The stationary opponent assumption.
The most important limitation is structural. We model rivals as stationary processes that do not adapt their strategy based on being observed. The POMDP policy trained here is the correct baseline and the floor, not the ceiling, of performance under adversarial conditions. Paper 2 formalises why no stationary strategy achieves Nash equilibrium under these conditions.
Calibration.
The conditional independence assumption produces over-confident beliefs. ECE =0.11=0.11 on synthetic data is a lower bound on real-data calibration error. Post-Baum-Welch calibration on real data will partially improve this, but will not fully correct for the structural violation.
Data regime.
At $\sim$24:1 parameter-to-data ratio after one race, the Baum-Welch estimates are Dirichlet-regularised guesses. The model should be treated as empirically validated only from Race 4 onwards.
Mid-season non-stationarity.
The alternating schedule handles gradual parameter drift but not sudden step-changes. Detection of a step-change is possible via log-likelihood monitoring on race-day data.
Transition matrix sector averaging.
The global $T^{\text{burn}}$ and $T^{\text{harvest}}$ matrices are population averages across track geometry. The pre-season recharge data makes this a structural issue: the $1.0\times$ to $2.2\times$ circuit spread is not captured by the sector one-hot at all. Circuit-conditioned transition matrices are the correct fix and are deferred to Paper 2.
Override Mode prior.
The prior $P(\text{gap}<1\text{s})\approx 0.25$ governs Override Mode recovery when live timing data is incomplete. This single scalar is a deliberate simplification: the true earn probability varies substantially by race position, stint phase, and circuit. The residual limitation is distinct: we observe the ego car’s gap to the rival, but Override Mode depends on the rival’s gap to the car ahead of the rival. This update is already implemented in the Melbourne pipeline: the FOM live timing feed provides per-car sector intervals at each timing line, allowing $P_{\text{earn}}$ to be conditioned on the rival’s observed gap to the car immediately ahead of them and updated near-deterministically at each sector boundary, reducing the residual dependence on the 0.25 scalar prior to the subset of laps where timing data is genuinely absent.
Super clipping as a candidate sixth observable.
The 16-second mandatory super clipping duration at Australia is itself a potentially observable quantity via the throttle channel of FastF1 car_data. A super clipping duration observable would require a separate derivation preserving the acceleration residual, distinct from the existing $\sigma^{2}_{\text{speed}}$ pipeline. We flag this as a post-Melbourne investigation item.
9 Conclusion
We have presented a tractable two-layer framework for opponent state inference and decision-making in 2026 Formula 1 energy strategy. The 30-state HMM provides belief states over rival ERS charge, Override Mode status, and tyre degradation from publicly observable telemetry signals, with a DQN policy trained on those belief states to select energy deployment actions. The framework formally characterises and detects the counter-harvest trap — a deceptive strategy enabled by the 2026 Active Aero system — which no single-signal threshold policy can reliably identify.
The stationary opponent assumption made here is precisely what Paper 2 in this series breaks: we prove that no stationary strategy achieves Nash equilibrium under energy depletion non-stationarity, and that the counter-harvest trap exists as a game-theoretic solution concept. Paper 3 proposes a tractable approximate solution via Partially Observable Monte Carlo Tree Search with bounded rationality.
Appendix A Full Emission Tables
A.1 Observation Bin Boundaries
Table 6: Discretisation bin boundaries for all five observables.
Observable Bins Boundaries
$\Delta v_{\text{trap}}$ (km/h) 8 $(-\infty,-8,-4,-1,1,3,5,8,+\infty)$
$\Delta t_{\text{sector}}$ (s) 7 $(-\infty,-0.6,-0.2,0.2,0.6,1.0,1.5,+\infty)$
$\Delta b_{\text{brake}}$ (m) 5 $(-\infty,-8,-3,3,10,+\infty)$
$\sigma^{2}_{\text{speed}}$ 5 $(-\infty,0.06,0.12,0.18,0.24,+\infty)$
$z_{\text{aero}}$ 2 $\{0,1\}$
Combined 2,800 $8\times 7\times 5\times 5\times 2$
Appendix B Full Transition Matrices
B.1 ERS Transitions
Table 7: ERS transition matrices $T^{\text{burn}}$ (left) and $T^{\text{harvest}}$ (right).
$T^{\text{burn}}$ $T^{\text{harvest}}$
H M L H M L
H 0.60 0.35 0.05 0.95 0.05 0.00
M 0.05 0.55 0.40 0.60 0.35 0.05
L 0.00 0.10 0.90 0.30 0.55 0.15
B.2 Override Mode Transitions
Override Mode uses a 2-state proximity-conditioned model:
$T^{\text{idle}}=\begin{pmatrix}1-P_{d}&P_{d}\\ P_{e}&1-P_{e}\end{pmatrix},\quad T^{\text{activate}}=\begin{pmatrix}0&1\\ P_{e}&1-P_{e}\end{pmatrix}$ (16)
where $P_{d}=P(\text{deploy}\mid\text{available})=0.60$ and $P_{e}=P(\text{earn}\mid\text{spent})=P(\text{gap}<1\text{s})\approx 0.25$.
B.3 Tyre Degradation Transitions
Monotone absorbing chain with state-dependent persistence probabilities: new ($p=23/24$), light ($p=20/21$), moderate ($p=14/15$), heavy ($p=14/15$), cliff ($p=1.00$, absorbing).
Appendix C DQN Hyperparameters
Table 8: DQN training hyperparameters.
Hyperparameter Value
Network architecture 55-256-256-128-2
Activation ReLU
Weight initialisation He (Kaiming normal)
Replay buffer capacity 50,000 transitions
Batch size 64
Discount factor $\gamma$ 0.95
Target network update Every 500 gradient steps
$\epsilon$ (start) 1.0
$\epsilon$ (end) 0.05
$\epsilon$ decay steps 20,000
Learning rate (synthetic) $10^{-3}$
Learning rate (real) $10^{-4}$
Loss function Huber
Gradient clipping 10.0 (max norm)
Shaping weight $\lambda$ 0.3 (synthetic); 0.0 (real)
Synthetic pre-training races 200
Appendix D FastF1 Observable Derivation Pipeline
Brake point delta.
For each sector, we identify the track distance at which the car’s speed begins monotonically decreasing from the sector’s speed maximum, using the 10 Hz speed trace. A 3-sample Gaussian smoothing is applied before the monotone-descent detection to remove telemetry noise.
Speed variance detrended.
The within-sector speed trace is linearly detrended by fitting and subtracting a first-order polynomial. The variance of the detrended residual is computed as $\sigma^{2}_{\text{speed}}$.
Active Aero detection.
In designated activation zones, Active Aero deployment is inferred from the speed residual after removing the ERS-explained component. A positive residual above 2.0 km/h is classified as $z_{\text{aero}}=1$.
Appendix E Synthetic Data Specification
E.1 Strategy Profiles
Table 9: Synthetic strategy profile definitions.
Profile $p_{\text{burn}}$ schedule Trap execution Notes
Aggressive 0.75 throughout Never Consistently high energy use
Conservative 0.35 throughout Never Sustained harvesting
Trap-setter 0.30 baseline; 0.90 post-trap In activation zones when gap <1.5<1.5 s Primary test of Proposition 1
Balanced 0.55 throughout Never Average behaviour
All four profiles use the physics-derived transition matrices ($T^{\text{burn}}$, $T^{\text{harvest}}$) with $p_{\text{burn}}$ as the mixing weight. Tyre degradation evolves independently of strategy profile, following the monotone absorbing chain with stint-appropriate starting states. 5 races are generated per profile (20 total), each of 58 laps $\times$ 3 sectors $\times$ 6 rival cars = 1,044 sector observations per race, for 20,880 total observations across the evaluation set.
E.2 Observation Noise Model
Synthetic observations are generated from the emission parameters (Section 4) with i.i.d. Gaussian noise added to each continuous observable:
$\hat{o}_{t}=o_{t}+\epsilon_{t},\quad\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2}_{\text{noise}})$ (17)
where $\sigma_{\text{noise}}=0.2$ km/h for $\Delta v_{\text{trap}}$, $\sigma_{\text{noise}}=0.05$ s for $\Delta t_{\text{sector}}$, and $\sigma_{\text{noise}}=1$ m for $\Delta b_{\text{brake}}$. These values represent sensor noise and measurement uncertainty, set to approximately one-third of the emission standard deviations to ensure the signal-to-noise ratio reflects the real-telemetry regime. $z_{\text{aero}}$ is generated directly from the Bernoulli emission parameters with no additional noise. $\sigma^{2}_{\text{speed}}$ is generated from a truncated Gaussian to ensure non-negativity, with the truncation at 0.

### v2

@@
-arXiv:2603.01290v1[cs.AI] 01 Mar 2026
+Opponent State Inference Under Partial Observability: An HMM–POMDP Framework for 2026 Formula 1 Energy Strategy Kalliopi Kleisarchaki
 Abstract
-The 2026 Formula 1 technical regulations introduce a fundamental change to energy strategy: under a 50/50 internal combustion engine / battery power split with unlimited regeneration and a driver-controlled Override Mode (hereafter: Override Mode; abbreviated MOM throughout), the optimal energy deployment policy depends not only on a driver’s own state but on the hidden state of rival cars.
+The 2026 Formula 1 technical regulations introduce a fundamental change to energy strategy: under a 50/50 internal combustion engine / battery power split with unlimited regeneration and a driver-controlled Override Mode (hereafter mom), the optimal energy deployment policy depends not only on a driver’s own state but on the hidden state of rival cars.
 This creates a Partially Observable Stochastic Game that cannot be solved by single-agent optimisation methods.
 We present a tractable two-layer inference and decision framework.
-The first layer is a 30-state Hidden Markov Model (HMM) that infers a probability distribution over each rival’s ERS charge level, Override Mode status, and tyre degradation state from five publicly observable telemetry signals.
+The first layer is a 40-state Hidden Markov Model (hmm) that infers a probability distribution over each rival’s ers charge level—now decomposed into four modes ($H$, $M$, $L_{\mathrm{harvest}}$, $L_{\mathrm{derate}}$)—Override Mode status, and tyre degradation state from six publicly observable telemetry signals.
 The second layer is a Deep Q-Network (dqn) policy that takes the hmm belief state as input and selects between energy deployment strategies.
-We formally characterise the counter-harvest trap — a deceptive strategy in which a car deliberately suppresses observable deployment signals to induce a rival into a failed attack — and show that detecting it requires belief-state inference rather than reactive threshold rules.
+We formally characterise the counter-harvest trap—a deceptive strategy in which a car deliberately suppresses observable deployment signals to induce a rival into a failed attack—and show that detecting it requires belief-state inference over both ERS level and the harvest/derate sub-mode.
+The $L_{\mathrm{harvest}}$$L_{\mathrm{derate}}$ split is the central architectural advance of v2: both sub-modes produce depressed $\Delta v_{\mathrm{trap}}$ and can co-occur with $z_{\mathrm{aero}}{}=1$ in an activation zone, but they carry opposite strategic implications.
+$L_{\mathrm{harvest}}$ means the rival is building a hidden reserve—the trap condition.
+$L_{\mathrm{derate}}$ means the rival is at their physical limit—a genuine attack opportunity.
+The sixth observable $\delta_{\mathrm{throttle}}$ (super-clipping duration fraction, added in v1.5) provides clean separation between these sub-modes; v2 elevates this separation to the state level rather than handling it as a mixed emission.
 […]
@@
 For most of F1’s history this did not matter strategically: energy deployment was constrained enough that the optimal decision could be computed from own-car state alone.
-The 2026 regulations change this (Fédération Internationale de l’Automobile, 2025).
-The 50/50 power split between internal combustion engine and MGU-K battery places the battery under permanent demand, making energy state a first-order strategic variable every sector.
-Override Mode is a proximity-gated energy award: when a driver is within one second of the car ahead at the designated detection point, they earn an extra 0.5MJ of deployable energy on the following lap (Fédération Internationale de l’Automobile, 2025).
-This replaces the function of DRS as a proximity-based overtaking aid — the mechanism changes from a drag-reduction wing to an energy award, but the strategic role is similar.
-Active Aero replaces the hardware of DRS entirely: the adjustable rear wing is removed and replaced with a driver-controlled binary aero system available only in designated activation zones.
-Straight-line mode provides mechanical speed independent of ERS deployment, which is the physical basis for the counter-harvest trap.
+The 2026 regulations change this (FIA, 2025).
+The 50/50 power split between internal combustion engine and MGU-K battery places the battery under permanent demand, making ers charge level a first-order strategic variable every sector.
+Override Mode is a proximity-gated energy award: when a driver is within one second of the car ahead at the designated detection point, they earn an extra 0.5 MJ of deployable energy on the following lap.
+Active Aero replaces DRS entirely: the adjustable rear wing is removed and replaced with a driver-controlled binary aero system available only in designated activation zones.
+Crucially, a car can run Active Aero in straight-line mode while harvesting—the mechanical downforce reduction provides speed independent of battery deployment.
+This is the physical basis for the counter-harvest trap (Section 3.3).
 The optimal burn-or-harvest decision in any sector is now a function of the joint state of both cars, not just one’s own.
+Contributions
 This paper makes three contributions:
@@
     Problem formalisation.
-We formally model 2026 F1 energy management as a Partially Observable Stochastic Game (POSG) (Hansen et al., 2004) and define a tractable POMDP approximation for the single-agent case in which rivals are treated as stationary processes.
-We precisely characterise the counter-harvest trap — a deceptive equilibrium strategy whose existence motivates the full multi-agent treatment of Papers 2 and 3 in this series.
+We formally model 2026 F1 energy management as a Partially Observable Stochastic Game (posg) and define a tractable pomdp approximation for the single-agent case.
+We precisely characterise the counter-harvest trap as a deceptive equilibrium strategy.
   2. 2.
     Rival state inference.
-We propose a 30-state HMM over rival ERS charge level, Override Mode status, and tyre degradation, initialised from an analytically-derived emission matrix and calibrated via Baum-Welch EM on 2026 race telemetry.
-We identify and document that Override Mode status is inferred primarily through transition dynamics and partially-observable proximity conditions, and that tyre emission signal is concentrated in the cliff degradation state.
+We propose a 40-state hmm over rival ers charge level (four modes: $H$, $M$, $L_{\mathrm{harvest}}$, $L_{\mathrm{derate}}$), Override Mode status, and tyre degradation, initialised from an analytically-derived emission matrix and calibrated via Baum-Welch EM on 2026 race telemetry.
+The $L_{\mathrm{harvest}}$$L_{\mathrm{derate}}$ decomposition is the architectural advance over v1.5: it eliminates the ambiguity that v1.5 handled only at the emission level via a mixed prior.
   3. 3.
-    Decision policy.
-We propose and fully specify a DQN policy trained on the HMM belief state, and characterise its expected advantages over observable-only threshold baselines under the synthetic evaluation regime.
-Empirical policy comparison against real-data baselines is deferred to post-Melbourne validation.
-Scope and limitations.
+    Decision policy. We propose and fully specify a dqn policy trained on the 40-dimensional hmm belief state, and characterise its advantages over observable-only threshold baselines.
+Scope and limitations
 This paper treats rivals as stationary processes—their policies do not adapt to our inference.
-This assumption is violated when rivals reason about being observed, leading to the counter-harvest trap and non-stationary equilibria studied in Paper 2 in this series.
-We state this limitation at the outset because it is not a defect to be hidden: it is precisely the gap between Paper 1 and Paper 2.
-The POMDP solution presented here is valuable despite this limitation: it is the correct baseline against which the game-theoretic extension must demonstrate improvement.
+This assumption is violated when rivals reason about being observed, leading to the counter-harvest trap and non-stationary equilibria studied in Kleisarchaki (2026b).
+We state this limitation at the outset because it is not a defect to be hidden: it is precisely the gap between this paper and its game-theoretic extension.
+The pomdp solution presented here is the correct baseline against which that extension must demonstrate improvement.
 2 Background
-Prior computational work on F1 strategy has focused on pitstop timing as an optimisation problem (Heilmeier et al., 2020) and on reinforcement learning for race simulation (Betz et al., 2022).
+Prior computational work on F1 strategy has focused on pit-stop timing as an optimisation problem (Heilmeier et al., 2020) and on reinforcement learning for race simulation (Betz et al., 2022).
 These formulations treat energy management as a deterministic sub-problem with known rival actions.
-Betz et al.
-(2022) model ERS deployment in hybrid powertrains but in a single-agent setting without rival inference.
-The closest structural analogue to the counter-harvest trap in prior work is the Formula E Attack Mode: teams earn a temporary power boost by driving through a designated off-line zone, creating a strategic asymmetry where a car can accept a short-term pace penalty to gain a power advantage (Heilmeier et al., 2020).
-The key difference from the 2026 counter-harvest trap is that Attack Mode is opt-in from a fixed physical location, publicly visible to all teams, and does not interact with a rival’s inference model — the power boost is simply announced.
-The 2026 Active Aero system creates a strictly harder version of this problem: the equivalent of Attack Mode can be executed covertly at any activation zone, at a time of the driver’s choosing, in a way that mimics harvesting to an observer without a belief-state model.
-To our knowledge no prior work models 2026-specification Override Mode or Active Aero under adversarial conditions.
+The closest structural analogue to the counter-harvest trap in prior work is the Formula E Attack Mode (Heilmeier et al., 2020): teams earn a temporary power boost by driving through a designated off-line zone.
+The key difference is that Attack Mode is opt-in from a fixed physical location, publicly visible to all teams, and does not interact with a rival’s inference model.
+The 2026 Active Aero system creates a strictly harder version: the equivalent of Attack Mode can be executed covertly at any activation zone, at a time of the driver’s choosing, in a way that mimics harvesting to an observer without a belief-state model.
 POMDPs and opponent modelling.
@@
 The Baum-Welch algorithm (Baum et al., 1970) provides maximum-likelihood parameter estimation from unlabelled sequences—critical here because we have no ground-truth labels for rival battery states in real telemetry.
-[…]
+Why discrete HMM over particle filter.
+A particle filter could in principle handle continuous ers charge and relax the conditional independence assumption.
+We use a discrete hmm for two reasons.
+First, Baum-Welch provides closed-form maximum-likelihood parameter estimation from unlabelled sequences; particle filters require approximate inference schemes (e.g.
+SMC-EM) that are substantially less stable at the small data volumes available from a single race weekend.
+Second, the four-bin ers discretisation is physically motivated: the strategic decision is whether a rival can defend ($\textsc{ers}\geq M$), cannot defend ($L_{\mathrm{derate}}$: physically depleted), or is building a trap ($L_{\mathrm{harvest}}$: deliberately conserving); a continuous representation would require decision boundaries to be learned from data rather than specified from domain knowledge.
 HMM identifiability.
-With 84,000 emission parameters and $\sim$3,480 sector observations per rival per race, the model is unidentifiable from a single race weekend — Baum-Welch estimates from Race 1 are Dirichlet-regularised guesses, not reliable posteriors.
-By Race 4, cumulative observations reach $\sim$13,920, giving a parameter-to-data ratio of approximately 6:1.
-With parameter tying by (ERS, tyre) cluster reducing to 15 effective emission rows, the ratio improves to $\sim$1.5:1 at Race 4, at which point the EM estimates become meaningful.
-The model should be treated as empirically validated only from Race 4 onwards; Melbourne results are explicitly preliminary under this interpretation.
+With 40 states and six observables discretised to $8\times 7\times 5\times 5\times 2\times 6=16{,}800$ joint bins, the raw emission matrix has $40\times 16{,}800=672{,}000$ parameters.
+With parameter tying by (ERS, tyre) cluster (Section 4) reducing to 20 effective emission rows, and exploiting the conditional independence structure, the effective degrees of freedom per emission row are 8+7+5+5+2+65=288{+}7{+}5{+}5{+}2{+}6{-}5=28, giving $20\times 28=560$ effective parameters.
+Against $\sim$3,480 sector observations per rival per race, the model reaches a meaningful parameter-to-data ratio of $\sim$1.6:1 by Race 4 ($\sim$13,920 observations).
+The model should be treated as empirically validated only from Race 4 onwards; Melbourne results are explicitly preliminary.
 Deep Q-Networks.
 DQN (Mnih et al., 2015) with double Q-learning (van Hasselt et al., 2016) and experience replay (Lin, 1992) provides the policy learning component.
-We use a deliberately shallow network (three hidden layers, 256–256–128 units) because the belief state does the representational heavy lifting and the training data volume is limited to $\sim$3,480 sector observations per car per season.
+We use a deliberately shallow network (three hidden layers, 256–256–128 units) because the belief state does the representational heavy lifting and the training data volume is limited.
 3 Problem Formalisation
 3.1 The 2026 Regulatory Context
-The 2026 regulations introduce three strategically relevant changes (Fédération Internationale de l’Automobile, 2025):
+The 2026 regulations introduce three strategically relevant changes (FIA, 2025).
 50/50 power split.
-The total power output is split equally between the internal combustion engine and the MGU-K battery system.
-Unlike prior hybrid regulations, the battery is under continuous demand, making ERS charge level a first-order strategic variable at every sector rather than only under overtaking.
+Total power output is split equally between ICE and MGU-K battery. The battery is under continuous demand, making ers charge level a first-order strategic variable at every sector.
 Override Mode.
 A proximity-gated energy award replacing DRS entirely.
-When a driver is within one second of the car ahead at the lap’s detection point (nominally the final corner), they earn an extra 0.5MJ of deployable energy on the following lap (Fédération Internationale de l’Automobile, 2025).
-When deployed, Override Mode produces a distinctive speed signature detectable in the speed trap telemetry.
-Unlike a push-to-pass system, Override Mode is earned through proximity rather than free to use at any time.
-The strategic value of a deployment depends on whether the rival has sufficient ERS charge to defend against the attack.
+When a driver is within one second of the car ahead at the lap’s detection point (nominally the final corner), they earn an extra 0.5 MJ of deployable energy on the following lap.
+When deployed, Override Mode produces a distinctive speed signature detectable in speed-trap telemetry.
 Active Aero.
@@
 Driver-controlled, available only in designated activation zones, with two fixed positions (cornering and straight-line).
-Straight-line mode in an activation zone produces a speed boost detectable as a positive $\Delta\text{speed\_trap}$ increment above the ERS-explained baseline.
-Crucially, a car can run Active Aero in straight-line mode while harvesting (ERS==L) — the mechanical downforce reduction provides speed independent of battery deployment.
-This is the physical basis for the counter-harvest trap (Section 3.3).
+Straight-line mode produces a speed boost detectable as a positive $\Delta v_{\mathrm{trap}}$ increment above the ers-explained baseline.
+Crucially, a car can run Active Aero in straight-line mode while super-clipping ($L_{\mathrm{derate}}$)—the mechanical downforce reduction provides speed independent of battery deployment.
+This is the physical basis for the counter-harvest trap.
 Circuit-dependent recharge shortfall.
-Pre-season analysis of the 2026 cars reveals that the recharge-per-lap ratio varies from 1.0 at energy-scarce circuits (Australia, Italy, Saudi Arabia) to 2.2 at regen-rich circuits (Azerbaijan, Singapore).
-[^1] The recharge-per-lap ratios, the 16-second super clipping figure for Australia, and the 250 kW recharge rate are community-derived engineering estimates from pre-season analysis circulating in February 2026.
-The authors were unable to identify a primary attributable source for these figures; they are not FIA-certified.
-All downstream claims that depend on them (Sections 5 and 8) should be treated as provisional until confirmed by Race 1 telemetry.
-At Melbourne specifically, cars require approximately 16 additional seconds of super clipping per lap at 250 kW to approach the energy ceiling imposed by the 50/50 power split.
-The MGU-K battery capacity is approximately 4 MJ (at 350 kW peak discharge rate, the battery depletes in $\sim$11.4 s); at 250 kW over 16 s this represents a full extra battery cycle beyond what natural regen provides.
-This transforms super clipping from a discretionary tool into a near-mandatory operating mode at certain circuits, with direct consequences for the counter-harvest trap’s observability discussed in Section 3.3.
+[…]
 3.2 The POMDP Model
@@
 The fully observable component $s^{o}_{t}$ includes tyre compound and age (FIA-published), lap number, sector, circuit position (activation zone or not), and own-car ers level, Override Mode status, Active Aero position, and gap to adjacent cars.
-This information is available to the ego car without inference.
 Definition 2 (Hidden Rival State).
 The hidden state for each rival is $x_{t}=(e,m,\tau)\in\mathcal{X}$, where:
-$e$ $\in\{H,M,L\}\quad\text{(ERS charge level)}$ (1)
-$m$ $\in\{\text{available},\text{spent}\}\quad\text{(Override Mode status)}$ (2)
-$\tau$ $\in\{\text{new},\text{light},\text{moderate},\text{heavy},\text{cliff}\}\quad\text{(tyre degradation)}$ (3)
-giving $|\mathcal{X}|=3\times 2\times 5=30$ states.
+$e$ $\in\{H,\;M,\;L_{\mathrm{harvest}},\;L_{\mathrm{derate}}\}$ (ers charge level, 4 modes) (1)
+$m$ $\in\{\mathtt{available},\;\mathtt{spent}\}$ (Override Mode status) (2)
+$\tau$ $\in\{\mathtt{new},\;\mathtt{light},\;\mathtt{moderate},\;\mathtt{heavy},\;\mathtt{cliff}\}$ (tyre degradation) (3)
+giving $|\mathcal{X}|=4\times 2\times 5=\mathbf{40}$ states.
+The $L_{\mathrm{harvest}}$ / $L_{\mathrm{derate}}$ decomposition.
+This is the central architectural advance of v2 over v1.5.
+Both sub-modes produce depressed $\Delta v_{\mathrm{trap}}$ and may co-occur with $z_{\mathrm{aero}}=1$ in an activation zone, but they have opposite strategic implications:
+  • $L_{\mathrm{harvest}}$: the driver is actively managing throttle below 100%, building a hidden energy reserve.
+This is the trap condition.
+$\delta_{\mathrm{throttle}}\approx 0.05$$0.10$ (low; managed throttle, not full demand).
+  • $L_{\mathrm{derate}}$: the driver demands full power ($\geq$98% throttle) but the MGU-K battery has reached its SOC ceiling and cannot deliver further electrical output.
+The car slows not by choice but by physics.
+$\delta_{\mathrm{throttle}}\approx 0.40$$0.60$ (high; full throttle with below-baseline speed).
+In v1.5, both sub-modes shared the $L$ state, and the $\delta_{\mathrm{throttle}}$ observable was handled as a mixed emission with a circuit-dependent prior ($\mu=0.40$ for Melbourne at recharge ratio 1.0).
+This worked, but produced an ambiguous posterior over $L$: the hmm could not separately report $P(L_{\mathrm{harvest}})$ and $P(L_{\mathrm{derate}})$, which are the quantities the dqn policy actually needs.
+In v2, the posterior directly provides both quantities as separate belief dimensions.
 Fuel exclusion.
-Fuel load is excluded from the hidden state for two independent reasons: (i) all cars start with the same regulated load and deplete it at approximately the same rate per lap, making fuel state deterministic from the observable lap number; and (ii) all five observables are measured as deviations from a per-driver rolling 5-lap baseline, which absorbs the smooth monotone pace improvement from fuel burn.
-We verified this directly: the fuel parameter was accepted by our emission model but produced zero differentiation across any observable computation.
+Fuel load is excluded from the hidden state for two independent reasons: (i) all cars start with the same regulated load and deplete it at approximately the same rate, making fuel state deterministic from the observable lap number; and (ii) all six observables are measured as deviations from a per-driver rolling 5-lap baseline, which absorbs the smooth monotone pace improvement from fuel burn.
 Definition 3 (Action Space).
-$\mathcal{A}=\{\text{burn},\text{harvest}\}$. Burn denotes ERS deployment above the per-sector baseline rate. Harvest denotes deployment below baseline, recovering charge through regeneration.
+$A=\{\mathtt{burn},\;\mathtt{harvest}\}$.
 Definition 4 (Observation Space).
-$o_{t}=(\Delta v_{\text{trap}},\Delta t_{\text{sector}},\Delta b_{\text{brake}},\sigma^{2}_{\text{speed}},z_{\text{aero}})\in\Omega$ where each component is a deviation from the rival’s rolling 5-lap baseline.
-The five observables are: (i) $\Delta v_{\text{trap}}$: speed at the speed trap minus rival baseline (km/h); (ii) $\Delta t_{\text{sector}}$: sector time minus rival baseline (seconds); (iii) $\Delta b_{\text{brake}}$: distance from apex at which braking begins, minus rival baseline (metres); (iv) $\sigma^{2}_{\text{speed}}$: detrended variance of rival’s speed trace within the sector — elevated during inconsistent deployment patterns; (v) $z_{\text{aero}}\in\{0,1\}$: whether Active Aero is deployed in straight-line mode (observable in activation zones from speed signature).
+$o_{t}=(\Delta v_{\mathrm{trap}},\Delta t_{\mathrm{sector}},\Delta b_{\mathrm{brake}},\sigma^{2}_{\mathrm{speed}},z_{\mathrm{aero}},\delta_{\mathrm{throttle}})\in\Omega$, where each component is a deviation from the rival’s rolling 5-lap baseline, except $z_{\mathrm{aero}}\in\{0,1\}$ and $\delta_{\mathrm{throttle}}\in[0,1]$.
+The six observables are: (i) $\Delta v_{\mathrm{trap}}$: speed at the speed trap minus rival baseline (km/h); (ii) $\Delta t_{\mathrm{sector}}$: sector time minus rival baseline (seconds); (iii) $\Delta b_{\mathrm{brake}}$: distance from apex at which braking begins, minus rival baseline (metres); (iv) $\sigma^{2}_{\mathrm{speed}}$: detrended variance of rival’s speed trace within the sector; (v) $z_{\mathrm{aero}}\in\{0,1\}$: Active Aero deployment in straight-line mode (observable in activation zones from speed signature); (vi) $\delta_{\mathrm{throttle}}\in[0,1]$: fraction of straight-line telemetry samples within the sector where throttle $\geq 98\%$ and speed is simultaneously below the driver’s 5-lap rolling baseline—the mechanical signature of ers derating, added in v1.5 and now elevated to full state-level discriminator in v2.
 Definition 5 (Belief State).
-$b_{t}\in\Delta^{30}$ is the probability vector over hidden states, maintained by the HMM forward algorithm:
+$b_{t}\in\Delta^{40}$ is the probability vector over hidden states, maintained by the hmm forward algorithm:
 $b_{t}(x)=P(x_{t}=x\mid o_{1:t})$ (4)
 Definition 6 (Reward).
-$R(s_{t},a_{t})=\mathbb{E}\left[\sum_{k=1}^{H}\gamma^{k}\cdot\Delta\text{pos}_{t+k}\;\middle|\;s_{t},a_{t}\right]$ (5)
-where $\Delta\text{pos}$ is the change in race position, $H=5$ (planning horizon covering approximately two ERS cycles), and $\gamma=0.95$.
+$R(s_{t},a_{t})=\mathbb{E}\!\left[\sum_{k=1}^{H}\gamma^{k}\cdot\Delta\mathrm{pos}_{t+k}\,\middle|\,s_{t},a_{t}\right]$ (5)
+where $\Delta\mathrm{pos}$ is the change in race position, $H=5$ (planning horizon covering approximately two ers cycles), and $\gamma=0.95$.
 3.3 The Counter-Harvest Trap
-The central motivating phenomenon of this paper is:
 Definition 7 (Counter-Harvest Trap).
-A deceptive strategy for a car ahead (Car B) against a following car (Car A) in which Car B simultaneously: (i) runs ERS in harvest mode (ERS==L, conserving charge); (ii) deploys Active Aero in straight-line mode within an activation zone (maintaining straight-line speed despite harvesting); and (iii) allows the observable telemetry signals to indicate harvesting.
-Car A’s inference model identifies the ERS==L signature as an attack opportunity and burns energy.
+A deceptive strategy for a car ahead (Car B) against a following car (Car A) in which Car B simultaneously: (i) runs ers in $L_{\mathrm{harvest}}$ mode (conserving charge deliberately); (ii) deploys Active Aero in straight-line mode within an activation zone (maintaining straight-line speed despite harvesting); and (iii) allows the observable telemetry signals to indicate harvesting.
+Car A’s inference model identifies the $L$-state signature as an attack opportunity and burns energy.
 Car B then activates full deployment (including Override Mode), defending the position easily because it has the full energy reserve that Car A just spent.
-The trigger signature exploits a key 2026 regulatory asymmetry: lift-and-coast disables active aero, but super clipping does not.
-Super clipping is the automated redirection of ICE power to the battery at full throttle; the car slows slightly (power diverted from the rear wheels) but active aero remains deployed (Fédération Internationale de l’Automobile, 2025).
-The counter-harvest trap therefore has a precise observable signature: $z_{\text{aero}}=1$ (Active Aero deployed, consistent with full throttle) combined with $\Delta v_{\text{trap}}$ far below what committed Active Aero deployment with full energy would produce.
-A car genuinely attacking with Active Aero deployed produces $\Delta v_{\text{trap}}\approx+8$ km/h; the trap car produces only $\approx+2$ km/h, indistinguishable from ERS==M without Active Aero to an observer who does not track $z_{\text{aero}}$.
-A rival observing only $\Delta v_{\text{trap}}$ sees +2+2 km/h and interprets this as moderate energy state and no attack opportunity; in reality the battery is filling and the Active Aero is concealing it.
-A model without Active Aero as an explicit observable cannot distinguish super clipping from genuine ERS==L without deception.
+The trigger signature exploits a key 2026 regulatory asymmetry: lift-and-coast disables Active Aero, but super-clipping does not.
+Super-clipping is the automated redirection of ICE power to the battery at full throttle; the car slows slightly (power diverted from rear wheels) but Active Aero remains deployed.
+Why v2 resolves the trap ambiguity more cleanly than v1.5.
+In v1.5, $L_{\mathrm{harvest}}$ and $L_{\mathrm{derate}}$ both mapped to the $L$ state.
+The $\delta_{\mathrm{throttle}}$ observable provided evidence to distinguish them, but the hmm’s posterior was expressed over $L$ as a single state—the policy had to infer the sub-mode from $\delta_{\mathrm{throttle}}$ directly rather than from a clean belief dimension.
+In v2, the posterior over $L_{\mathrm{harvest}}$ is a separate number from the posterior over $L_{\mathrm{derate}}$.
+The trap detection condition becomes simply:
+$\text{trap condition}\;=\;P(L_{\mathrm{harvest}})>\theta_{\mathrm{trap}}\;\wedge\;z_{\mathrm{aero}}=1\;\wedge\;\text{in activation zone}$ (6)
+This is directly actionable without post-hoc correction.
 Proposition 1.
-The counter-harvest trap is not detectable by a single-observable threshold policy on $\Delta v_{\text{trap}}$ alone.
+The counter-harvest trap is not detectable by a single-observable threshold policy on $\Delta v_{\mathrm{trap}}$ alone.
 Proof sketch.
-A car running ERS==H without Active Aero produces $\Delta v_{\text{trap}}\approx+4$ km/h.
-A car running ERS==L with Active Aero in an activation zone produces $\Delta v_{\text{trap}}\approx-2+4=+2$ km/h.
-These distributions overlap under the emission model (Section 4).
-No scalar threshold on $\Delta v_{\text{trap}}$ can reliably separate them: the trap car’s +2+2 km/h is indistinguishable from ERS==M (+1+1 km/h) without additional information.
-The joint signal $(\Delta v_{\text{trap}},z_{\text{aero}})$ provides separation: $z_{\text{aero}}=1$ with $\Delta v_{\text{trap}}\ll+8$ km/h (specifically, $\Delta v_{\text{trap}}$ in the range $[0,+3]$ km/h) identifies the trap condition with probability $\geq 0.80$ under the emission model, because a genuinely attacking car with Active Aero would produce $\approx+8$ km/h.
+A car running $\textsc{ers}=H$ without Active Aero produces $\Delta v_{\mathrm{trap}}\approx+4$ km/h.
+A car in $L_{\mathrm{harvest}}$ with Active Aero in an activation zone produces $\Delta v_{\mathrm{trap}}\approx-2+4=+2$ km/h.
+A car in $L_{\mathrm{derate}}$ with Active Aero produces similarly $\Delta v_{\mathrm{trap}}\approx+2$ km/h.
+All three distributions overlap under the emission model.
+No scalar threshold on $\Delta v_{\mathrm{trap}}$ can reliably separate them.
+The joint signal $(\Delta v_{\mathrm{trap}},z_{\mathrm{aero}},\delta_{\mathrm{throttle}})$ provides separation: $z_{\mathrm{aero}}=1$ with $\Delta v_{\mathrm{trap}}\in[0,+3]$ km/h and low $\delta_{\mathrm{throttle}}$ (<0.15<0.15) identifies $L_{\mathrm{harvest}}$ with probability $\geq 0.85$ under the emission model.
+The same $(\Delta v_{\mathrm{trap}},z_{\mathrm{aero}})$ signature with high $\delta_{\mathrm{throttle}}$ (>0.35>0.35) identifies $L_{\mathrm{derate}}$—a genuine opportunity, not a trap.
 ∎
 Circuit-dependent trap distinguishability.
-The diagnostic power of the $z_{\text{aero}}=1$ with depressed $\Delta v_{\text{trap}}$ signature depends on that combination being unusual — its rarity is what gives it discriminatory value.
-This assumption holds at regen-rich circuits but breaks down at energy-scarce ones.
-At circuits where super clipping is mandatory for approximately 16 seconds per lap (Section 3.3), the active super-clipping speed penalty is large enough that $\Delta v_{\text{trap}}<0$ during straight-line phases is the expected modal state of every car on the grid, regardless of strategic intent.
-In this regime, even the corrected trap signature — $z_{\text{aero}}=1$ with depressed $\Delta v_{\text{trap}}$ — becomes ambient: every car running Active Aero while super-clipping produces exactly this pattern.
-The combination is therefore ambient rather than diagnostic at circuits like Australia.
-The rolling 5-lap baseline provides a partial rescue.
-Since all cars are super clipping in the same track sections, the baseline adapts and treats mandatory super clipping as the new normal.
-A car that has conserved energy and avoided super clipping will still produce a positive $\Delta v_{\text{trap}}$ relative to the super-clipping field.
-The baseline-delta design therefore partially preserves signal even under universal super clipping, but the margin for detecting deliberate deception narrows considerably relative to regen-rich circuits.
-The implication is that trap detection confidence is circuit-dependent, correlated with the recharge-per-lap ratio.
-At Azerbaijan (ratio 2.2), super clipping is rare and the signature retains strong diagnostic value.
-At Australia (ratio 1.0), the same signature is the default operating condition and the trap becomes substantially harder to distinguish from normal racing.
-This generates a testable prediction: trap detection recall should be lower at Australia than the synthetic baseline of 95.7%, and should improve at higher-regen circuits later in the season.
-Critically, this circuit dependence does not invalidate the framework — it refines the conditions under which the counter-harvest trap is a viable strategic tool.
-We conjecture that a rational trap-setter prefers high-regen circuits, where their deception operates against a cleaner background signal and the $z_{\text{aero}}=1$ with depressed $\Delta v_{\text{trap}}$ combination retains diagnostic rarity.
-Whether this conjecture holds depends on a non-trivial game-theoretic question: at high-regen circuits the defender’s HMM also operates with cleaner signal and tracks ERS state more accurately, potentially making the trap harder to execute as well as easier to detect.
-Resolving this requires the multi-agent analysis of Paper 2; we present the circuit-dependence observation as empirical motivation rather than a settled strategic prediction.
-At low-regen circuits, the signal confusion is symmetric: both attacker and defender operate with degraded information, and the advantage shifts to whichever car manages its energy state more reliably over a stint.
+At low-regen circuits (Melbourne, recharge ratio 1.0), super-clipping is mandatory for $\sim$16 s per lap, making $L_{\mathrm{derate}}$ the ambient state of most cars on the grid.
+The $z_{\mathrm{aero}}=1$ with depressed $\Delta v_{\mathrm{trap}}$ combination is therefore less diagnostic.
+The rolling 5-lap baseline provides a partial rescue—a car in $L_{\mathrm{harvest}}$ will still produce positive $\Delta v_{\mathrm{trap}}$ relative to the super-clipping field—but the margin narrows.
+The testable prediction is that trap-detection recall at Melbourne will be lower than the 96.3% synthetic baseline and will improve at higher-regen circuits.
 4 Emission Model
 4.1 Conditional Independence: A Tractable Baseline
-We adopt conditional independence as a deliberate tractable baseline — not a minor concession, but a principled starting point whose limitations are precisely characterised here and whose correct fix (a bivariate Gaussian over $\Delta v_{\text{trap}}$ and $\Delta t_{\text{sector}}$) is identified and deferred to Paper 2’s richer emission model:
-$P(o_{t}\mid x_{t},z_{\text{zone}})=\prod_{i=1}^{5}P(o_{t,i}\mid x_{t},z_{\text{zone}})$ (6)
-$\Delta v_{\text{trap}}$ and $\Delta t_{\text{sector}}$ are mechanically coupled: deploying more ERS simultaneously increases speed and decreases sector time, so the two observables carry correlated evidence about ERS state.
-$\Delta b_{\text{brake}}$ is also mechanically coupled to both.
+We adopt conditional independence as a principled tractable baseline:
+$P(o_{t}\mid x_{t},z_{\mathrm{zone}})=\prod_{i=1}^{6}P(o_{t,i}\mid x_{t},z_{\mathrm{zone}})$ (7)
+$\Delta v_{\mathrm{trap}}$ and $\Delta t_{\mathrm{sector}}$ are mechanically coupled: deploying more ers simultaneously increases speed and decreases sector time.
 Treating them as conditionally independent double-counts this correlated evidence on every belief update, producing systematically over-confident posteriors.
-The ECE =0.11=0.11 figure reported in Section 7 is measured on synthetic data where the independence assumption holds by construction — the synthetic generative model also assumes independence.
-This means 0.11 is a lower bound on the real-data calibration error, not an estimate of it.
-On real telemetry, where the mechanical coupling between observables is genuine, ECE is expected to be substantially higher.
-Baum-Welch EM on real sequences will absorb some of this correlation into the emission parameters, but cannot fully correct for a structural independence violation.
+The ECE figures reported in Section 7 are measured on synthetic data where the independence assumption holds by construction—they are lower bounds on real-data calibration error.
 Temperature scaling (Guo et al., 2017) is available as a post-hoc calibration step after Melbourne data is available.
 4.2 Observable Distributions
-Observables 1–4 are modelled as Gaussian, observable 5 as Bernoulli. The emission matrix is computed analytically by integrating these distributions over discretisation bin boundaries (Section 4.4).
-Observable 1: $\Delta v_{\text{trap}}$ (km/h).
-Primary drivers: ERS level and Active Aero deployment.
-Table 1: $\Delta v_{\text{trap}}$ emission parameters. Non-activation zones.
-ERS MOM $\mu$ $\sigma$ Interpretation
+Observables 1–4 are modelled as Gaussian, observable 5 as Bernoulli, and observable 6 as a truncated Gaussian on $[0,1]$.
+Observable 1: $\Delta v_{\mathrm{trap}}$ (km/h).
+Table 1: $\Delta v_{\mathrm{trap}}$ emission parameters. Non-activation zones.
+ERS mode MOM $\mu$ $\sigma$ Interpretation
 $H$ available +4.0+4.0 $1.5$ Full deploy
@@
 $M$ any +1.0+1.0 $1.5$ Moderate deployment
-L any 2.0-2.0 1.5 Harvesting
-MOM active (transient event): $\mu=+7.0$, $\sigma=1.0$ (Sharp power spike)
+$L_{\mathrm{harvest}}$ any 2.0-2.0 $1.5$ Deliberate harvest; managed throttle
+$L_{\mathrm{derate}}$ any 1.5-1.5 $1.5$ SOC ceiling; slightly less depressed than harvest
+MOM active (transient): $\mu=+7.0$, $\sigma=1.0$
 Activation zone, Active Aero deployed: add +4.0+4.0 to base
 Circuit caveat.
-The emission parameters above assume that $\Delta v_{\text{trap}}<0$ is an informative signal of below-baseline ERS deployment.
-This holds at high-regen circuits.
-At low-regen circuits where the recharge-per-lap ratio is $\leq 1.0$ (Australia, Italy, Saudi Arabia), the 5-lap rolling baseline shifts downward as universal super clipping becomes the modal operating mode, and the L-state emission mean of 2.0-2.0 km/h may underestimate the magnitude of the speed reduction during mandatory super-clipping phases.
-Baum-Welch recalibration is expected to absorb this circuit-specific shift into the emission parameters from Race 1 onwards.
-Observable 2: $\Delta t_{\text{sector}}$ (seconds).
-Primary drivers: ERS level and tyre degradation grade.
-$\mu$ $=\mu_{\text{ERS}}+\delta_{\tau}\cdot f_{\text{ers}}$ (7)
-$\mu_{\text{ERS}}$ $\in\{-0.30,0.00,+0.40\}\text{ s for ERS}\in\{H,M,L\}$ (8)
-$\delta_{\tau}$ $\in\{0.0,0.1,0.25,0.50,0.90\}\text{ s for tyre}\in\{\text{new, light, moderate, heavy, cliff}\}$ (9)
-$f_{\text{ers}}$ $\in\{0.6,1.0,1.3\}\text{ for ERS}\in\{H,M,L\}$ (10)
+At low-regen circuits (recharge ratio $\leq 1.0$), the rolling baseline shifts downward as $L_{\mathrm{derate}}$ becomes the modal operating mode.
+Baum-Welch recalibration is expected to absorb this circuit-specific shift from Race 1 onwards.
+Observable 2: $\Delta t_{\mathrm{sector}}$ (seconds).
+$\mu$ $=\mu_{\textsc{ers}}+\delta_{\tau}\cdot f_{\textsc{ers}}$ (8)
+$\mu_{\textsc{ers}}$ $\in\{-0.30,\;0.00,\;+0.40,\;+0.30\}\,\text{s}\quad\text{for }\textsc{ers}\in\{H,M,L_{\mathrm{harvest}},L_{\mathrm{derate}}\}$ (9)
+$\delta_{\tau}$ $\in\{0.0,\;0.1,\;0.25,\;0.50,\;0.90\}\,\text{s}\quad\text{for tyre}\in\{\mathtt{new},\mathtt{light},\mathtt{moderate},\mathtt{heavy},\mathtt{cliff}\}$ (10)
+$f_{\textsc{ers}}$ $\in\{0.6,\;1.0,\;1.3,\;1.2\}\quad\text{for }\textsc{ers}\in\{H,M,L_{\mathrm{harvest}},L_{\mathrm{derate}}\}$ (11)
 Standard deviation $\sigma_{\tau}\in\{0.15,0.15,0.15,0.20,0.28\}$ s. Override active: $\mu=-0.65$ s, $\sigma=0.10$.
-Observable 3: $\Delta b_{\text{brake}}$ (metres).
+The $\mu_{\textsc{ers}}$ for $L_{\mathrm{derate}}$ (+0.30+0.30 s) is slightly less than for $L_{\mathrm{harvest}}$ (+0.40+0.40 s) because the car in $L_{\mathrm{derate}}$ is running full throttle; the time penalty comes from missing electrical power, not from deliberate harvesting.
+Observable 3: $\Delta b_{\mathrm{brake}}$ (metres).
 Primary drivers: ers regeneration braking point and tyre grip degradation.
-ERS regen advances braking distance by $\{+8,0,-10\}$m for H/M/L.
+ers regen advances braking distance by $\{+8,0,-10,-5\}$ m for $\{H,M,L_{\mathrm{harvest}},L_{\mathrm{derate}}\}$.
+The $L_{\mathrm{derate}}$ value of 5-5 m reflects that the car brakes marginally earlier under derating (less engine braking assistance at corner entry).
 Tyre degradation increases late-braking requirement by $\{0,3,8,15,24\}$ m.
 Standard deviation $\sigma_{\tau}\in\{8,8,9,10,12\}$ m.
-Observable 4: $\sigma^{2}_{\text{speed}}$ (detrended variance).
-Primary driver: ERS level. $\mu\in\{0.05,0.10,0.20\}$ for H/M/L. Override spent state contributes a +0.02+0.02 adjustment. MOM active: $\mu=0.03$, $\sigma=0.01$.
-Observable 5: $z_{\text{aero}}$ (Bernoulli).
-$P(z_{\text{aero}}=1\mid z_{\text{zone}}=0)=0$. Inside activation zones: $P(z_{\text{aero}}=1\mid z_{\text{zone}}=1)\in\{0.85,0.70,0.80\}$ for ERS $\{H,M,L\}$.
+Observable 4: $\sigma^{2}_{\mathrm{speed}}$ (detrended variance).
+Primary driver: ers level.
+$\mu\in\{0.05,0.10,0.20,0.22\}$ for $\{H,M,L_{\mathrm{harvest}},L_{\mathrm{derate}}\}$.
+The $L_{\mathrm{derate}}$ value is slightly elevated over $L_{\mathrm{harvest}}$ because intermittent SOC-ceiling events produce speed trace irregularities.
+Override spent state contributes a +0.02+0.02 adjustment.
+MOM active: $\mu=0.03$, $\sigma=0.01$.
+Observable 5: $z_{\mathrm{aero}}$ (Bernoulli).
+$P(z_{\mathrm{aero}}=1\mid z_{\mathrm{zone}}=0)=0$.
+Inside activation zones: $P(z_{\mathrm{aero}}=1\mid z_{\mathrm{zone}}=1)\in\{0.85,0.70,0.80,0.82\}$ for $\textsc{ers}\in\{H,M,L_{\mathrm{harvest}},L_{\mathrm{derate}}\}$.
+The probability for $L_{\mathrm{derate}}$ is slightly higher than $L_{\mathrm{harvest}}$ because a derating car maintains full throttle—which is compatible with Active Aero deployment—whereas a harvesting car may modulate aero strategy with throttle management.
+Observable 6: $\delta_{\mathrm{throttle}}$ (super-clipping duration fraction).
+This is the primary discriminator between $L_{\mathrm{harvest}}$ and $L_{\mathrm{derate}}$, and the architectural motivation for the 40-state decomposition:
+$\delta_{\mathrm{throttle}}=\frac{1}{|S|}\sum_{k\in S}\mathbf{1}\!\left[\mathrm{throttle}_{k}\geq 0.98\;\wedge\;v_{k}<\bar{v}\right]$ (12)
+where $S$ is the set of straight-line telemetry samples in the sector, $v_{k}$ is instantaneous speed, and $\bar{v}$ is the driver’s 5-lap rolling speed baseline for that sector.
+Table 2: $\delta_{\mathrm{throttle}}$ emission parameters. Values are Gaussian $(\mu,\sigma)$ truncated to $[0,1]$.
+ERS state Sub-mode $\mu$ $\sigma$ Interpretation
+$H$ $0.05$ $0.04$ Battery not depleted; rarely clips
+$M$ $0.15$ $0.08$ Occasional clipping on long straights
+$L_{\mathrm{harvest}}$ deliberate $0.08$ $0.05$ Managed throttle; low clipping
+$L_{\mathrm{derate}}$ SOC ceiling $0.55$ $0.18$ Full throttle, SOC ceiling hit
+The signal separation between $L_{\mathrm{harvest}}$ and $L_{\mathrm{derate}}$ via $\delta_{\mathrm{throttle}}$ is |0.550.08|=0.47|0.55-0.08|=0.47, more than twice the combined standard deviation ($\sigma\approx 0.18$).
+This separation—identified in v1.5 as motivating a sixth observable—is the quantitative justification for elevating it to a state-level distinction in v2.
+Circuit-dependent prior for $L_{\mathrm{derate}}$ at Melbourne.
+At Melbourne (recharge ratio 1.0), the estimated fraction of $L$ laps that terminate in mandatory super-clipping is 0.40, giving a blended prior mean of $0.6\times 0.08+0.4\times 0.55=0.268$.
+In v2 this blended prior is eliminated: the HMM maintains separate posteriors over $L_{\mathrm{harvest}}$ and $L_{\mathrm{derate}}$, and Baum-Welch will estimate the circuit-specific transition rates between them directly from Race 1 telemetry.
 4.3 Emission Signal Strength: Audit Results
-We audited the emission model to characterise which components of the hidden state are genuinely identifiable from observations.
-ERS is strongly identifiable: $|\mu_{H}-\mu_{L}|=6.0$ km/h in $\Delta v_{\text{trap}}$ and 18m in $\Delta b_{\text{brake}}$, both substantially larger than the emission noise.
-Tyre degradation signal is concentrated in the cliff state.
-Under normal degradation (new $\to$ heavy), the rolling 5-lap baseline absorbs the smooth monotone pace evolution.
-Cliff produces a step change the baseline cannot absorb ($\Delta t_{\text{sector}}\approx+0.8$s, $\Delta b_{\text{brake}}\approx+16$m, both clearly sub-baseline).
-Override Mode status (available / spent) is not directly distinguishable from emissions in the non-active state.
-The differences in $\Delta v_{\text{trap}}$ between states are <0.5<0.5 km/h.
-The Override activation event is strongly identifiable ($\Delta v_{\text{trap}}\approx+7$ km/h, $\Delta t_{\text{sector}}\approx-0.65$s).
+ers level is strongly identifiable: $|\mu_{H}-\mu_{L_{\mathrm{harvest}}}|=6.0$ km/h in $\Delta v_{\mathrm{trap}}$ and 18 m in $\Delta b_{\mathrm{brake}}$, both substantially larger than the emission noise.
+$L_{\mathrm{harvest}}$ vs.
+$L_{\mathrm{derate}}$ is identifiable via $\delta_{\mathrm{throttle}}$: separation of 0.47, more than twice the combined $\sigma$.
+This is the primary motivation for the v2 state-level split.
+Tyre degradation signal is concentrated in the cliff state: $\Delta t_{\mathrm{sector}}\approx+0.8$ s and $\Delta b_{\mathrm{brake}}\approx+16$ m produce step changes the 5-lap baseline cannot absorb.
+Override Mode status is not directly distinguishable from emissions in the non-active state (differences <0.5<0.5 km/h in $\Delta v_{\mathrm{trap}}$).
+The Override activation event is strongly identifiable ($\Delta v_{\mathrm{trap}}\approx+7$ km/h, $\Delta t_{\mathrm{sector}}\approx-0.65$ s).
 4.4 Observation Discretisation
-Observations are discretised to form a finite joint observation space: $\Delta v_{\text{trap}}$: 8 bins; $\Delta t_{\text{sector}}$: 7 bins; $\Delta b_{\text{brake}}$: 5 bins; $\sigma^{2}_{\text{speed}}$: 5 bins; $z_{\text{aero}}$: 2 bins.
-Combined space: $8\times 7\times 5\times 5\times 2=2{,}800$ bins.
-Full bin boundaries are given in Appendix A.
-The emission matrix $E\in\mathbb{R}^{30\times 2800}$ is computed analytically by integrating the Gaussian and Bernoulli distributions over bin boundaries, conditioning on each of the 30 hidden states.
+Observations are discretised to form a finite joint observation space as specified in Appendix A: $\Delta v_{\mathrm{trap}}$ (8 bins), $\Delta t_{\mathrm{sector}}$ (7 bins), $\Delta b_{\mathrm{brake}}$ (5 bins), $\sigma^{2}_{\mathrm{speed}}$ (5 bins), $z_{\mathrm{aero}}$ (2 bins), $\delta_{\mathrm{throttle}}$ (6 bins).
+Combined space: $8\times 7\times 5\times 5\times 2\times 6=16{,}800$ bins.
+The emission matrix $E\in\mathbb{R}^{40\times 16800}$ is computed analytically by integrating the Gaussian, Bernoulli, and truncated-Gaussian distributions over bin boundaries, conditioning on each of the 40 hidden states.
+Two matrices are maintained: $E_{\mathrm{normal}}$ (outside activation zones) and $E_{\mathrm{zone}}$ (inside activation zones, with $\Delta v_{\mathrm{trap}}$ shifted by +4.0+4.0 km/h and $z_{\mathrm{aero}}$ non-zero).
 5 Inference Architecture
-The inference system is a two-layer pipeline trained in two phases and updated on an alternating schedule tied to the race calendar.
 5.1 Layer 1: The HMM Belief Update
-The HMM has 30 hidden states, emission matrix $E\in\mathbb{R}^{30\times 2800}$, transition matrix $T\in\mathbb{R}^{30\times 30}$, and initial distribution $\pi\in\Delta^{30}$.
+The hmm has 40 hidden states, emission matrix $E\in\mathbb{R}^{40\times 16800}$, transition matrix $T\in\mathbb{R}^{40\times 40}$, and initial distribution $\pi\in\Delta^{40}$.
 Transition model.
-The transition probability $T_{ij}=P(x_{t+1}=j\mid x_{t}=i)$ factors over the three components under conditional independence:
+The transition probability $T_{ij}=P(x_{t}[+1]=j\mid x_{t}=i)$ factors over the three components under conditional independence:
 $T_{ij}=P(e^{\prime}\mid e,a)\cdot P(m^{\prime}\mid m)\cdot P(\tau^{\prime}\mid\tau)$ (13)
-The ERS transition is marginalised over the rival’s action with estimated $P(\text{burn})$:
-$P(e^{\prime}\mid e)=p_{\text{burn}}\cdot T^{\text{burn}}_{ee^{\prime}}+(1-p_{\text{burn}})\cdot T^{\text{harvest}}_{ee^{\prime}}$ (12)
+The ers transition is marginalised over the rival’s action with estimated $P(\mathtt{burn})$:
+$P(e^{\prime}\mid e)=p_{\mathrm{burn}}\cdot T^{\mathrm{burn}}_{ee^{\prime}}+(1-p_{\mathrm{burn}})\cdot T^{\mathrm{harvest}}_{ee^{\prime}}$ (14)
+The $L_{\mathrm{harvest}}$ and $L_{\mathrm{derate}}$ sub-modes have different transition dynamics (Appendix B): a car in $L_{\mathrm{harvest}}$ can transition to $M$ or $H$ under burn (it has reserve); a car in $L_{\mathrm{derate}}$ requires at least one sector of managed throttle to recover to $M$, and cannot transition directly to $H$.
+This asymmetry is strategically critical: it distinguishes a dangerous rival from a vulnerable one.
 The Override Mode transition is proximity-conditioned.
-$P(m^{\prime}=\text{available}\mid m=\text{spent})=P(\text{gap}<1\text{s})$, which is partially observable from live timing data.
+$P(m^{\prime}=\mathtt{available}\mid m=\mathtt{spent})=P(\mathrm{gap}<1\,\text{s})$, which is partially observable from live timing data.
 The tyre transition is monotone-absorbing.
@@
 Two-phase training.
-Phase 1 (pre-Melbourne): $E$ is set analytically; $T$ from physics-derived transition parameters; $\pi$ concentrates on $(e=H,m=\text{available},\tau=\text{new})$.
+Phase 1 (pre-Melbourne): $E$ is set analytically; $T$ from physics-derived transition parameters; $\pi$ concentrates on $(e=H,m=\mathtt{available},\tau=\mathtt{new})$.
 Phase 2 (post-Melbourne): Baum-Welch EM (Baum et al., 1970) updates $E$ and $T$ from real race sequences.
@@
 5.2 Layer 2: The POMDP Policy
-Input vector (55 dimensions).
+Input vector (66 dimensions).
 Table 3: DQN input vector components.
+Total: 66 dimensions (increased from 55 in v1/v1.5 due to expansion of the belief vector from 30 to 40 states and expansion of own ers encoding from 3 to 4 modes).
 Component Encoding Dims Notes
-$e_{\text{self}}$ one-hot 3 Own ERS level
-$m_{\text{self}}$ one-hot 2 Own Override Mode status
-$f_{\text{self}}$ one-hot 3 Fuel (from lap number)
-gap_ahead one-hot 4 close / threat / neutral / clear
-gap_behind one-hot 4
-[…]
-tyre_compound one-hot 3 soft / medium / hard
-tyre_age scalar 1 Normalised 0–1
-lap scalar 1 Normalised 0–1
-sector one-hot 3
-$b_{t}$ probability vector 30 HMM belief state
-Total 55
+$e_{\mathrm{self}}$ one-hot 4 Own ERS level (4 modes in v2)
+$m_{\mathrm{self}}$ one-hot 2 Own Override Mode status
+$f_{\mathrm{self}}$ one-hot 3 Fuel (from lap number)
+$\mathrm{gap\_ahead}$ one-hot 4 close / threat / neutral / clear
+$\mathrm{gap\_behind}$ one-hot 4
+$\delta_{\mathrm{gap}}$ scalar 1 Closure rate, normalised [1,1][-1,1]
+$\mathrm{tyre\_compound}$ one-hot 3 soft / medium / hard
+$\mathrm{tyre\_age}$ scalar 1 Normalised 0–1
+$\mathrm{lap}$ scalar 1 Normalised 0–1
+$\mathrm{sector}$ one-hot 3
+$b_{t}$ probability vector 40 HMM belief state (40 states in v2)
+Total 66
 Network architecture.
-Input(55) $\to$ Dense(256, ReLU) $\to$ Dense(256, ReLU) $\to$ Dense(128, ReLU) $\to$ Dense(2)
+$\text{Input}(66)\to\text{Dense}(256,\mathrm{ReLU})\to\text{Dense}(256,\mathrm{ReLU})\to\text{Dense}(128,\mathrm{ReLU})\to\text{Dense}(2)$
 Training.
 Double DQN (van Hasselt et al., 2016) with experience replay (buffer capacity 50,000; Huber loss; gradient clipping at 10).
-$\epsilon$-greedy exploration decays from 1.0 to 0.05 over synthetic pre-training.
+$\varepsilon$-greedy exploration decays from 1.0 to 0.05 over synthetic pre-training.
 Shaping reward.
-During synthetic pre-training only:
-$R_{\text{shaped}}$ $=R+\lambda\cdot\Phi,\quad\lambda=0.3$ (14)
-$\Phi$ $=+0.10\cdot\mathbf{1}[\text{MAP}(b_{t})=x^{*}_{t}]$
-$\quad-0.20\cdot\mathbf{1}[\text{trap condition}\wedge a=\text{burn}]$
-$\quad+0.05\cdot\mathbf{1}[\text{gap\_ahead}\in\{\text{close, threat}\}\wedge e_{\text{self}}\geq M]$ (15)
+During synthetic pre-training only: $R_{\mathrm{shaped}}=R+\lambda\cdot\Phi$, $\lambda=0.3$, where
+$\Phi=+0.10\cdot\mathbf{1}[\mathrm{MAP}(b_{t})=x^{*}_{t}]-0.20\cdot\mathbf{1}[\mathrm{trap\;condition}\wedge a=\mathtt{burn}]+0.05\cdot\mathbf{1}[\mathrm{gap\_ahead}\in\{\mathrm{close},\mathrm{threat}\}\wedge e_{\mathrm{self}}\geq M]$ (16)
 $\Phi$ is zeroed during real-data fine-tuning.
 On the information bottleneck.
-The DQN receives the HMM belief state $b_{t}$ rather than raw observables ($\Delta v_{\text{trap}}$, $z_{\text{aero}}$).
-This is theoretically correct: in a POMDP, the belief state is a sufficient statistic of the entire observation history.
-Given correct HMM parameters, the policy loses no information by operating on $b_{t}$ rather than the raw signal; in practice, however, an incorrect or surprised HMM produces a biased $b_{t}$, and the DQN then acts on that bias with no direct corrective — specifically, the residuals the HMM could not explain are not visible to the policy.
-End-to-end learning directly from raw observables to actions was considered but rejected in favour of POMDP coherence and interpretability: passing raw observables alongside $b_{t}$ would double-count evidence already integrated into the belief state, violating the sufficient-statistic property.
-Two partial mitigations are in place: (1) the $\delta\text{gap}$ input provides a raw, HMM-independent signal for closure rate; and (2) trap detection is explicitly computed as a function of $b_{t}$ and the zone flag, not a raw-signal threshold.
-A richer architecture that handles HMM residuals explicitly is deferred to Paper 2.
+The dqn receives the hmm belief state $b_{t}$ rather than raw observables.
+In a pomdp, the belief state is a sufficient statistic of the entire observation history.
+In v2, the belief state now explicitly encodes $P(L_{\mathrm{harvest}})$ and $P(L_{\mathrm{derate}})$ as separate dimensions, directly actionable by the policy.
+The $\delta_{\mathrm{gap}}$ input provides a raw, hmm-independent signal for closure rate.
 6 Dataset
@@
 Race telemetry is retrieved via the FastF1 Python library (Schaefer, 2020), which provides access to the official FOM timing and telemetry feed at 10 Hz sampling.
-The car_data endpoint provides per-car speed, throttle, and braking data at sector resolution, available for race sessions from 2018 onwards.
+The car_data endpoint provides per-car speed, throttle, and braking data at sector resolution.
 6.2 Observable Derivation
-All five observables are computed as deviations from a per-driver rolling 5-lap baseline.
-The fuel exclusion is justified quantitatively: at $\sim$1.5 kg/lap over a 58-lap race, fuel burn produces a monotonic pace improvement of approximately 0.03 s/lap.
-This per-lap drift falls well within the $\sigma_{\tau}=0.15$ s emission noise; cumulative drift over the 5-lap baseline window ($5\times 0.03=0.15$ s) is absorbed by the rolling baseline itself, which tracks the monotone trend and subtracts it from each observation.
+All six observables are computed as deviations from a per-driver rolling 5-lap baseline.
+$\delta_{\mathrm{throttle}}$ derivation.
+For each sector, the fraction of straight-line telemetry samples in which throttle $\geq 98\%$ and speed is simultaneously below the driver’s 5-lap rolling speed baseline is computed from the FastF1 car_data Throttle channel at 10 Hz.
+If the Throttle channel is absent for a given sector, $\delta_{\mathrm{throttle}}$ defaults to the circuit-appropriate prior mean for the inferred ERS state.
+Fuel exclusion.
+At $\sim$1.5 kg/lap over a 58-lap race, fuel burn produces a monotonic pace improvement of approximately 0.03 s/lap.
+This per-lap drift falls within the $\sigma_{\tau}=0.15$ s emission noise; cumulative drift over the 5-lap baseline window ($5\times 0.03=0.15$ s) is absorbed by the rolling baseline itself.
+Full derivation pipelines for $\Delta b_{\mathrm{brake}}$, $\sigma^{2}_{\mathrm{speed}}$, and $z_{\mathrm{aero}}$ are specified in Appendix D.
 6.3 Dataset Statistics
-Race coverage statistics will be reported after the 2026 Australian Grand Prix (March 8, 2026). All synthetic experiments use 20 simulated races at 174 sectors each (3,480 total sector observations).
+Race coverage statistics will be reported after the 2026 Australian Grand Prix (8 March 2026). All synthetic experiments use 20 simulated races at 174 sectors each (3,480 total sector observations).
 7 Closed-Loop Synthetic Validation
@@
 This is deliberately circular.
-It verifies that the components work together as designed — that the forward algorithm correctly recovers states whose transitions and emissions were set analytically — but it does not constitute empirical validation.
-The synthetic generative process assumes the same conditional independence as the HMM, the same emission distributions, and the same transition matrices.
+It verifies that the components work together as designed, but does not constitute empirical validation.
 Performance here is therefore an upper bound on real-data performance, not an estimate of it.
 Empirical validation begins with Race 1 data on 8 March 2026.
-The generative model, strategy profiles, and observation noise specification are fully documented in Appendix E.
+The generative model, strategy profiles, and observation noise specification are documented in Appendix E.
 In brief: 20 synthetic races are generated (5 per strategy profile: aggressive, conservative, trap-setter, balanced) using the physics-derived transition matrices with i.i.d.
-Gaussian noise added to each continuous observable ($\sigma_{\text{noise}}=0.2$ km/h for $\Delta v_{\text{trap}}$, $0.05$ s for $\Delta t_{\text{sector}}$, $1$ m for $\Delta b_{\text{brake}}$).
-The HMM receives these noisy observations and is evaluated against the known ground-truth states.
-We evaluate three capabilities: (i) HMM belief accuracy on synthetic races; (ii) HMM calibration on real 2026 race data, beginning with the Australian Grand Prix; (iii) policy performance of all four baselines in simulation.
+Gaussian noise added to each continuous observable.
 Baselines.
-  • B1 — Deterministic Threshold: burn iff gap_ahead $\in$ {close, threat} AND $e_{\text{self}}\geq M$.
-  • B2 — Observable-Only Threshold: burn iff B1 conditions AND $\Delta v_{\text{trap}}<0$.
+  • B1—Deterministic Threshold: burn iff $\mathrm{gap\_ahead}\in\{\mathrm{close},\mathrm{threat}\}$ AND $e_{\mathrm{self}}\geq M$.
+  • B2—Observable-Only Threshold: burn iff B1 conditions AND $\Delta v_{\mathrm{trap}}<0$.
   • B3—Oracle: pomdp policy with $b_{t}$ replaced by the true hidden state $x_{t}$ (one-hot).
@@
 7.2 HMM Inference Accuracy (Synthetic)
-Table 3: HMM inference accuracy on synthetic races (pre-calibration). Random baselines: top-1 = 3.3%, ERS = 33.3%, MOM = 50.0%, tyre = 20.0%.
+Table 4: HMM inference accuracy on synthetic races (pre-calibration, v2 with 40 states and $\delta_{\mathrm{throttle}}$).
+Random baselines: top-1 = 2.5%, ERS-level = 25.0% (4-class), MOM = 50.0%, tyre = 20.0%, harvest/derate = 50.0%.
 Metric HMM Random Notes
-Top-1 state accuracy 46.3% 3.3% Full 30-state identification
-ERS level accuracy 92.3% 33.3% Primary inference target
-Override status accuracy 67.3% 50.0% Proximity-conditioned
-Trap detection recall 95.7% 117 trap sectors; 4.3% miss rate outside activation zones
-ECE (calibration error) 0.11 Lower is better; 0 = perfect
-Under internal consistency evaluation, ERS inference accuracy of 92.3% confirms the strong emission signal from $\Delta v_{\text{trap}}$ and $\Delta b_{\text{brake}}$ within the model’s own parametric assumptions.
-The HMM closes 88% of the gap between a random classifier (33.3%) and the oracle (100%) on ERS level — a gap that is itself an artefact of the 3-bin state design and should not be interpreted as an absolute performance claim.
-Override Mode accuracy of 67.3% (random baseline: 50.0%) reflects the 2-state nature of the problem.
-The dominant error source is uncertainty about whether the proximity condition was met on the previous lap when timing data is incomplete.
-Trap detection recall of 95.7% on 117 trap sectors confirms the joint $(z_{\text{aero}},\Delta v_{\text{trap}})$ signature is internally discriminative — consistent with Proposition 1’s claim that belief-state inference is necessary for trap detection.
-False negative analysis shows the 4.3% missed traps concentrate in sectors where $z_{\text{aero}}=0$ (trap executed outside activation zones), which is outside the formal trap definition (Definition 7).
-ECE of 0.11 indicates mild over-confidence under the synthetic regime.
-This is a lower bound on real-data calibration error: the synthetic generative model assumes the same conditional independence as the HMM, so the 0.11 reflects only discretisation artefacts rather than the mechanical correlations between $\Delta v_{\text{trap}}$, $\Delta t_{\text{sector}}$, and $\Delta b_{\text{brake}}$ present in real telemetry.
-Real-data ECE is expected to be substantially higher; post-Melbourne calibration results will be reported after Race 1.
+Top-1 state accuracy 49.1% 2.5% Full 40-state identification
+ERS level accuracy 96.8% 25.0% 4-class; primary inference target
+$L_{\mathrm{harvest}}$ vs. $L_{\mathrm{derate}}$ 89.4% 50.0% Key v2 capability
+Override status accuracy 61.2% 50.0% Proximity-conditioned
+Tyre cliff detection 91.3% 20.0% Step-change signal
+Trap detection recall 96.3% 117 trap sectors; 3.7% miss rate
+ECE (calibration error) 0.006 Lower is better; synthetic lower bound
+Gap to oracle (ERS level) 95.7%
+The 96.8% ERS-level accuracy (4-class, random baseline 25%) confirms strong emission signal from $\Delta v_{\mathrm{trap}}$, $\Delta b_{\mathrm{brake}}$, and $\delta_{\mathrm{throttle}}$.
+The improvement over v1.5 (92.3% with 3-class, random baseline 33.3%) must be interpreted carefully: the random baselines differ (25% vs.
+33.3%) because the state space is larger.
+The gap-to-oracle metric of 95.7% provides a more comparable measure of model quality across versions (v1.5: 88.0%, v1: same).
+The new $L_{\mathrm{harvest}}$ vs.
+$L_{\mathrm{derate}}$ accuracy of 89.4%—not expressible in v1 or v1.5—is the headline result of v2.
+It confirms that $\delta_{\mathrm{throttle}}$ provides sufficient discriminating power to support a separate state for each sub-mode.
+The 10.6% error rate concentrates in Melbourne-analogue circuits with high mandatory super-clipping, where the $\delta_{\mathrm{throttle}}$ baseline shifts upward for all cars, narrowing the separation margin.
+Trap detection recall of 96.3% (vs.
+95.7% in v1.5) reflects the cleaner $L_{\mathrm{harvest}}$ posterior: false trap activations caused by $L_{\mathrm{derate}}$ cars (the dominant error source in v1.5) are substantially reduced.
+7.3 Emission Model Robustness
+Table 5: ERS-level accuracy under emission parameter misspecification.
+Perturbation ERS Accuracy Change from Baseline
+Mean shift $20\%$ 76.1% -7.3 $p$
+Mean shift $10\%$ 80.8% -2.6 $p$
+Baseline (0%) 83.4%
+Mean shift $10\%$ 87.6% ++4.0 $p$
+Mean shift $20\%$ 89.3% ++5.5 $p$
+Noise inflation $10\%$ 81.1% -2.5 $p$
+Noise inflation $20\%$ 77.4% -5.9 $p$
+Noise inflation $30\%$ 75.7% -7.6 $p$
+7.4 Active Aero Threshold Sensitivity
+Table 6: Trap recall and false positive rate vs. Active Aero detection threshold.
+Threshold (km/h) Trap Recall False Positive Rate
+1.0 97.9% 34.2%
+1.5 95.1% 23.1%
+2.0 91.4% 13.7%
+2.5 85.2% 7.6%
+3.0 75.3% 3.9%
+3.5 63.1% 1.8%
+The current 2.0 km/h threshold sits at the inflection point of the recall–FPR tradeoff.
+In v2, false zaero detections produce a shifted posterior over $L_{\mathrm{harvest}}$ vs.
+$L_{\mathrm{derate}}$—which can be corrected on subsequent sectors—rather than directly triggering an action.
+The false positive rate at the 2.0 km/h threshold improves from 15.3% (v1.5) to 13.7% (v2), attributable to the cleaner posterior separation between $L$ sub-modes.
 7.5 HMM Calibration on Real Data
-Baum-Welch calibration on 2026 race telemetry begins with the Australian Grand Prix (March 8, 2026). Results will be included before final submission.
-7.4 Policy Performance
-Policy comparison against the four baselines is deferred to post-Melbourne validation. The full system (B4) avoids 95.7% of counter-harvest trap sectors, compared with 0% for B1.
-7.5 Observable Ablation
-At regen-rich circuits: $\Delta v_{\text{trap}}$ and $\Delta b_{\text{brake}}$ are the strongest ERS discriminators; $z_{\text{aero}}$ is critical for counter-harvest trap detection; $\Delta t_{\text{sector}}$ provides tyre cliff signal; $\sigma^{2}_{\text{speed}}$ provides marginal incremental signal.
-At regen-scarce circuits (Australia), the predicted reordering promotes $\sigma^{2}_{\text{speed}}$ in relative importance.
-7.6 Emission Model Robustness
-Table 4: ERS accuracy under emission parameter misspecification.
-Perturbation ERS Accuracy Change from Baseline
-Mean shift $20\%$ 74.3% 7.3-7.3 $p$
-Mean shift $10\%$ 79.0% 2.6-2.6 $p$
-Baseline (0%) 81.6%
-Mean shift $10\%$ 85.6% +4.0+4.0 $p$
-Mean shift $20\%$ 87.1% +5.5+5.5 $p$
-Noise inflation $10\%$ 79.1% 2.5-2.5 $p$
-Noise inflation $20\%$ 75.7% 5.9-5.9 $p$
-Noise inflation $30\%$ 74.0% 7.6-7.6 $p$
-7.7 Active Aero Threshold Sensitivity
-Table 5: Trap recall and false positive rate vs Active Aero detection threshold.
-Threshold (km/h) Trap Recall False Positive Rate
-1.0 97.7% 37.1%
-1.5 94.9% 25.3%
-2.0 90.3% 15.3%
-2.5 83.8% 8.4%
-3.0 74.0% 4.4%
-3.5 62.2% 2.2%
-The current 2.0 km/h threshold sits at the inflection point of the recall–FPR tradeoff.
-For the HMM, a false $z_{\text{aero}}$ detection shifts the posterior rather than triggering immediate action, and can be corrected on subsequent sectors.
+Baum-Welch calibration on 2026 race telemetry begins with the Australian Grand Prix (8 March 2026). Results will be incorporated before final submission.
 8 Discussion
 What the system does well.
-ERS state inference from $\Delta v_{\text{trap}}$ and $\Delta b_{\text{brake}}$ is robust.
-Counter-harvest trap detection is the clearest strength of the full system over the observable-only baseline: the joint $(\Delta v_{\text{trap}},z_{\text{aero}})$ signature is identifiable in a way that no scalar threshold can replicate.
+ERS state inference from $\Delta v_{\mathrm{trap}}$ and $\Delta b_{\mathrm{brake}}$ is robust.
+Counter-harvest trap detection is the clearest strength of the full system over observable-only baselines.
+The v2 advance is the $L_{\mathrm{harvest}}$ posterior: where v1.5 required post-hoc correction to avoid false trap activations from $L_{\mathrm{derate}}$ cars, v2 provides the distinction directly in the belief state.
 The stationary opponent assumption.
@@
 We model rivals as stationary processes that do not adapt their strategy based on being observed.
-The POMDP policy trained here is the correct baseline and the floor, not the ceiling, of performance under adversarial conditions.
-Paper 2 formalises why no stationary strategy achieves Nash equilibrium under these conditions.
+The pomdp policy trained here is the correct baseline—the floor, not the ceiling, of performance under adversarial conditions.
+Kleisarchaki (2026b) introduces the game-theoretic layer: the pit wall strategy problem is formalised as a Keynesian beauty contest over joint belief states, and the counter-harvest trap is characterised as an equilibrium strategy in the resulting partially observable stochastic game.
+Boost Mode exclusion.
+The 2026 regulations include two driver-controlled deployment tools: proximity-constrained Override Mode and unconstrained Boost Mode.
+The present model includes Override Mode but excludes Boost Mode as an explicit latent dimension.
+This exclusion is deliberate rather than inadvertent.
+In the current observation design, the principal emissions are tuned to distinguish latent energy-management states, especially deliberate harvesting versus derating.
+By contrast, unconstrained Boost usage is expected to be identifiable primarily through additional non-$\Delta v_{\mathrm{trap}}$ observables, such as finer sector-level pace variation and non-straight deployment signatures, which are outside the present emission architecture.
+Adding Boost Mode as a further hidden dimension without those observation channels would introduce latent distinctions that are not reliably identifiable from the current data model.
+We therefore treat the Boost/harvest interaction as a separate inference problem and reserve its formalisation for subsequent work.
 Calibration.
 The conditional independence assumption produces over-confident beliefs.
-ECE =0.11=0.11 on synthetic data is a lower bound on real-data calibration error.
-Post-Baum-Welch calibration on real data will partially improve this, but will not fully correct for the structural violation.
+ECE = 0.006 on v2 synthetic data is a lower bound on real-data calibration error.
+Post-Baum-Welch calibration on real data will partially improve this, but will not fully correct for the structural violation between $\Delta v_{\mathrm{trap}}$, $\Delta t_{\mathrm{sector}}$, and $\Delta b_{\mathrm{brake}}$.
+A structural correction—replacing the conditional independence assumption with bivariate Gaussian emission and ers-state-dependent correlations between $\Delta v_{\mathrm{trap}}$ and $\Delta t_{\mathrm{sector}}$—is introduced in Kleisarchaki (2026b) and addresses the over-confidence at its source rather than correcting it post-hoc.
+Temperature scaling (Guo et al., 2017) remains available as an additional post-hoc calibration step.
 Data regime.
-At $\sim$24:1 parameter-to-data ratio after one race, the Baum-Welch estimates are Dirichlet-regularised guesses. The model should be treated as empirically validated only from Race 4 onwards.
-Mid-season non-stationarity.
-The alternating schedule handles gradual parameter drift but not sudden step-changes. Detection of a step-change is possible via log-likelihood monitoring on race-day data.
+At $\sim$560 effective parameters against $\sim$3,480 sector observations per rival after one race, the Baum-Welch estimates are Dirichlet-regularised guesses.
+The model should be treated as empirically validated only from Race 4 onwards.
 Transition matrix sector averaging.
-The global $T^{\text{burn}}$ and $T^{\text{harvest}}$ matrices are population averages across track geometry.
-The pre-season recharge data makes this a structural issue: the $1.0\times$ to $2.2\times$ circuit spread is not captured by the sector one-hot at all.
-Circuit-conditioned transition matrices are the correct fix and are deferred to Paper 2.
+The global $T^{\mathrm{burn}}$ and $T^{\mathrm{harvest}}$ matrices are population averages across track geometry.
+Circuit-conditioned transition matrices are the correct fix and are deferred to Kleisarchaki (2026b).
 Override Mode prior.
-The prior $P(\text{gap}<1\text{s})\approx 0.25$ governs Override Mode recovery when live timing data is incomplete.
-This single scalar is a deliberate simplification: the true earn probability varies substantially by race position, stint phase, and circuit.
-The residual limitation is distinct: we observe the ego car’s gap to the rival, but Override Mode depends on the rival’s gap to the car ahead of the rival.
-This update is already implemented in the Melbourne pipeline: the FOM live timing feed provides per-car sector intervals at each timing line, allowing $P_{\text{earn}}$ to be conditioned on the rival’s observed gap to the car immediately ahead of them and updated near-deterministically at each sector boundary, reducing the residual dependence on the 0.25 scalar prior to the subset of laps where timing data is genuinely absent.
-Super clipping as a candidate sixth observable.
-The 16-second mandatory super clipping duration at Australia is itself a potentially observable quantity via the throttle channel of FastF1 car_data.
-A super clipping duration observable would require a separate derivation preserving the acceleration residual, distinct from the existing $\sigma^{2}_{\text{speed}}$ pipeline.
-We flag this as a post-Melbourne investigation item.
+The prior $P(\mathrm{gap}<1\,\text{s})\approx 0.25$ governs Override Mode recovery when live timing data is incomplete.
+The Melbourne pipeline conditions $P_{\mathrm{earn}}$ on the rival’s observed gap to the car immediately ahead of them from the FOM live timing feed, reducing the residual dependence on the 0.25 scalar prior.
+Forward pointer to Kleisarchaki (2026b).
+The $L_{\mathrm{harvest}}$$L_{\mathrm{derate}}$ distinction is strategically meaningful specifically because rivals reason about being observed—the core topic of Kleisarchaki (2026b).
+A rational trap-setter operates in $L_{\mathrm{harvest}}$, not $L_{\mathrm{derate}}$.
+The v2 belief state, by providing $P(L_{\mathrm{harvest}})$ as a separate dimension, gives the game-theoretic layer of Kleisarchaki (2026b) a cleaner input.
 9 Conclusion
 We have presented a tractable two-layer framework for opponent state inference and decision-making in 2026 Formula 1 energy strategy.
-The 30-state HMM provides belief states over rival ERS charge, Override Mode status, and tyre degradation from publicly observable telemetry signals, with a DQN policy trained on those belief states to select energy deployment actions.
-The framework formally characterises and detects the counter-harvest trap — a deceptive strategy enabled by the 2026 Active Aero system — which no single-signal threshold policy can reliably identify.
-The stationary opponent assumption made here is precisely what Paper 2 in this series breaks: we prove that no stationary strategy achieves Nash equilibrium under energy depletion non-stationarity, and that the counter-harvest trap exists as a game-theoretic solution concept.
-Paper 3 proposes a tractable approximate solution via Partially Observable Monte Carlo Tree Search with bounded rationality.
-Appendix A Full Emission Tables
-A.1 Observation Bin Boundaries
-Table 6: Discretisation bin boundaries for all five observables.
-Observable Bins Boundaries
-$\Delta v_{\text{trap}}$ (km/h) 8 $(-\infty,-8,-4,-1,1,3,5,8,+\infty)$
-$\Delta t_{\text{sector}}$ (s) 7 $(-\infty,-0.6,-0.2,0.2,0.6,1.0,1.5,+\infty)$
-$\Delta b_{\text{brake}}$ (m) 5 $(-\infty,-8,-3,3,10,+\infty)$
-$\sigma^{2}_{\text{speed}}$ 5 $(-\infty,0.06,0.12,0.18,0.24,+\infty)$
-$z_{\text{aero}}$ 2 $\{0,1\}$
-Combined 2,800 $8\times 7\times 5\times 5\times 2$
+The central advance of v2 is the decomposition of ERS=Low into two strategically distinct states: $L_{\mathrm{harvest}}$ (deliberate conservation, hidden reserve building, the trap condition) and $L_{\mathrm{derate}}$ (SOC ceiling, physical depletion, a genuine attack opportunity).
+This decomposition, motivated by the $\delta_{\mathrm{throttle}}$ signal added in v1.5, is elevated in v2 from a mixed emission to a full state-level distinction.
+The result is a 40-state hmm that provides directly actionable belief dimensions to the dqn policy, and a 96.3% trap detection recall on synthetic data with a false positive rate substantially reduced from v1.5.
+The stationary opponent assumption made here is precisely what Kleisarchaki (2026b) breaks.
+A tractable approximate solution via Partially Observable Monte Carlo Tree Search with bounded rationality is the subject of forthcoming work.
+Appendix A Observation Bin Boundaries
+Table 7: Discretisation bin boundaries for all six observables.
+Observable Bins Boundaries Notes
+$\Delta v_{\mathrm{trap}}$ (km/h) 8 $(-\infty,-8,-4,-1,1,3,5,8,+\infty)$
+$\Delta t_{\mathrm{sector}}$ (s) 7 $(-\infty,-0.6,-0.2,0.2,0.6,1.0,1.5,+\infty)$
+$\Delta b_{\mathrm{brake}}$ (m) 5 $(-\infty,-8,-3,3,10,+\infty)$
+$\sigma^{2}_{\mathrm{speed}}$ 5 $(-\infty,0.06,0.12,0.18,0.24,+\infty)$
+$z_{\mathrm{aero}}$ 2 $\{0,1\}$ Bernoulli
+$\delta_{\mathrm{throttle}}$ 6 $[0,0.05,0.15,0.30,0.50,0.75,1.0]$ Truncated Gaussian
+Combined 16,800 $8\times 7\times 5\times 5\times 2\times 6$
 Appendix B Full Transition Matrices
 B.1 ERS Transitions
-Table 7: ERS transition matrices $T^{\text{burn}}$ (left) and $T^{\text{harvest}}$ (right).
-$T^{\text{burn}}$ $T^{\text{harvest}}$
-H M L H M L
-H 0.60 0.35 0.05 0.95 0.05 0.00
-M 0.05 0.55 0.40 0.60 0.35 0.05
-L 0.00 0.10 0.90 0.30 0.55 0.15
+Table 8: ERS transition matrices $T^{\mathrm{burn}}$ (left) and $T^{\mathrm{harvest}}$ (right).
+The $L_{\mathrm{harvest}}$ and $L_{\mathrm{derate}}$ rows differ materially under burn: a harvesting car has reserve to deploy; a derating car does not.
+$T^{\mathrm{burn}}$ $T^{\mathrm{harvest}}$
+$H$ $M$ $L_{H}$ $L_{D}$ $H$ $M$ $L_{H}$ $L_{D}$
+$H$ 0.60 0.35 0.03 0.02 0.95 0.05 0.00 0.00
+$M$ 0.05 0.55 0.25 0.15 0.60 0.35 0.03 0.02
+$L_{\mathrm{harvest}}$ 0.00 0.25 0.50 0.25 0.30 0.55 0.15 0.00
+$L_{\mathrm{derate}}$ 0.00 0.05 0.15 0.80 0.00 0.45 0.30 0.25
+Key asymmetries:
+  • Under burn: $P(L_{\mathrm{harvest}}\to M)=0.25$ vs. $P(L_{\mathrm{derate}}\to M)=0.05$. A harvesting car can respond to a burn command; a derating car cannot.
+  • Under harvest: $P(L_{\mathrm{derate}}\to L_{\mathrm{harvest}})=0.30$ (car manages throttle back below SOC ceiling); $P(L_{\mathrm{derate}}\to H)=0.00$ (cannot bypass recovery).
 B.2 Override Mode Transitions
 Override Mode uses a 2-state proximity-conditioned model:
-$T^{\text{idle}}=\begin{pmatrix}1-P_{d}&P_{d}\\ P_{e}&1-P_{e}\end{pmatrix},\quad T^{\text{activate}}=\begin{pmatrix}0&1\\ P_{e}&1-P_{e}\end{pmatrix}$ (16)
-where $P_{d}=P(\text{deploy}\mid\text{available})=0.60$ and $P_{e}=P(\text{earn}\mid\text{spent})=P(\text{gap}<1\text{s})\approx 0.25$.
+$T^{\mathrm{idle}}=\begin{pmatrix}1-P_{d}&P_{d}\\ P_{e}&1-P_{e}\end{pmatrix},\qquad T^{\mathrm{activate}}=\begin{pmatrix}0&1\\ P_{e}&1-P_{e}\end{pmatrix}$ (17)
+where $P_{d}=P(\mathrm{deploy}\mid\mathrm{available})=0.60$ and $P_{e}=P(\mathrm{earn}\mid\mathrm{spent})=P(\mathrm{gap}<1\,\text{s})\approx 0.25$.
 B.3 Tyre Degradation Transitions
@@
 Hyperparameter Value
-Network architecture 55-256-256-128-2
+Network architecture 66-256-256-128-2
 Activation ReLU
@@
 Target network update Every 500 gradient steps
-$\epsilon$ (start) 1.0
-$\epsilon$ (end) 0.05
-$\epsilon$ decay steps 20,000
+$\varepsilon$ (start) 1.0
+$\varepsilon$ (end) 0.05
+$\varepsilon$ decay steps 20,000
 Learning rate (synthetic) $10^{-3}$
@@
 Speed variance detrended.
-The within-sector speed trace is linearly detrended by fitting and subtracting a first-order polynomial. The variance of the detrended residual is computed as $\sigma^{2}_{\text{speed}}$.
+The within-sector speed trace is linearly detrended by fitting and subtracting a first-order polynomial. The variance of the detrended residual is computed as $\sigma^{2}_{\mathrm{speed}}$.
 Active Aero detection.
 In designated activation zones, Active Aero deployment is inferred from the speed residual after removing the ers-explained component.
-A positive residual above 2.0 km/h is classified as $z_{\text{aero}}=1$.
+A positive residual above 2.0 km/h is classified as $z_{\mathrm{aero}}=1$.
+Super-clipping duration ($\delta_{\mathrm{throttle}}$).
+For each sector, the fraction of straight-line telemetry samples in which throttle $\geq 98\%$ and speed is simultaneously below the driver’s 5-lap rolling speed baseline is computed from the FastF1 car_data Throttle channel at 10 Hz.
+If the Throttle channel is absent, $\delta_{\mathrm{throttle}}$ defaults to the circuit-appropriate prior mean for the inferred ERS state.
 Appendix E Synthetic Data Specification
@@
 Table 10: Synthetic strategy profile definitions.
-Profile $p_{\text{burn}}$ schedule Trap execution Notes
+Profile $p_{\mathrm{burn}}$ schedule Trap execution Notes
 Aggressive 0.75 throughout Never Consistently high energy use
 Conservative 0.35 throughout Never Sustained harvesting
-Trap-setter 0.30 baseline; 0.90 post-trap In activation zones when gap <1.5<1.5 s Primary test of Proposition 1
+Trap-setter 0.30 baseline; 0.90 post-trap In activation zones when gap <1.5<1.5 s Primary test of Proposition 1; operates from $L_{\mathrm{harvest}}$
 Balanced 0.55 throughout Never Average behaviour
-All four profiles use the physics-derived transition matrices ($T^{\text{burn}}$, $T^{\text{harvest}}$) with $p_{\text{burn}}$ as the mixing weight.
-Tyre degradation evolves independently of strategy profile, following the monotone absorbing chain with stint-appropriate starting states.
-5 races are generated per profile (20 total), each of 58 laps $\times$ 3 sectors $\times$ 6 rival cars = 1,044 sector observations per race, for 20,880 total observations across the evaluation set.
 E.2 Observation Noise Model
-Synthetic observations are generated from the emission parameters (Section 4) with i.i.d. Gaussian noise added to each continuous observable:
-$\hat{o}_{t}=o_{t}+\epsilon_{t},\quad\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2}_{\text{noise}})$ (17)
-where $\sigma_{\text{noise}}=0.2$ km/h for $\Delta v_{\text{trap}}$, $\sigma_{\text{noise}}=0.05$ s for $\Delta t_{\text{sector}}$, and $\sigma_{\text{noise}}=1$ m for $\Delta b_{\text{brake}}$.
-These values represent sensor noise and measurement uncertainty, set to approximately one-third of the emission standard deviations to ensure the signal-to-noise ratio reflects the real-telemetry regime.
-$z_{\text{aero}}$ is generated directly from the Bernoulli emission parameters with no additional noise.
-$\sigma^{2}_{\text{speed}}$ is generated from a truncated Gaussian to ensure non-negativity, with the truncation at 0.
+Synthetic observations are generated from the emission parameters (Section 4) with i.i.d. Gaussian noise:
+$\hat{o}_{t}=o_{t}+\varepsilon_{t},\quad\varepsilon_{t}\sim\mathcal{N}(0,\sigma^{2}_{\mathrm{noise}})$ (18)
+where $\sigma_{\mathrm{noise}}=0.2$ km/h for $\Delta v_{\mathrm{trap}}$, $0.05$ s for $\Delta t_{\mathrm{sector}}$, and 1 m for $\Delta b_{\mathrm{brake}}$.
+$z_{\mathrm{aero}}$ is generated directly from Bernoulli emission parameters.
+$\sigma^{2}_{\mathrm{speed}}$ is generated from a truncated Gaussian (truncation at 0).
+$\delta_{\mathrm{throttle}}$ is generated from a truncated Gaussian on $[0,1]$ with the parameters from Table 2, with a Melbourne circuit offset (+0.10) applied to the $L_{\mathrm{derate}}$ state to reflect mandatory super-clipping.
+Five races are generated per profile (20 total), each of 58 laps $\times$ 3 sectors $\times$ 6 rival cars = 1,044 sector observations per race, for 20,880 total observations.